### v2

# GePUP-ES: High-order Energy-stable Projection Methods for Incompressible Navier-Stokes Equations with No-slip Conditions

Abstract
Inspired by the unconstrained PPE (UPPE) formulation [Liu, Liu, & Pego 2007 Comm. Pure Appl. Math., 60 pp. 1443], we previously proposed the GePUP formulation [Zhang 2016 J. Sci. Comput., 67 pp. 1134] for numerically solving incompressible Navier-Stokes equations (INSE) on no-slip domains. In this paper, we propose GePUP-E and GePUP-ES, variants of GePUP that feature (a) electric boundary conditions with no explicit enforcement of the no-penetration condition, (b) equivalence to the no-slip INSE, (c) exponential decay of the divergence of an initially non-solenoidal velocity, and (d) monotonic decrease of the kinetic energy. Different from UPPE, the GePUP-E and GePUP-ES formulations are of strong forms and are designed for finite volume/difference methods under the framework of method of lines. Furthermore, we develop semi-discrete algorithms that preserve (c) and (d) and fully discrete algorithms that are fourth-order accurate for velocity both in time and in space. These algorithms employ algebraically stable time integrators in a black-box manner and only consist of solving a sequence of linear equations in each time step. Results of numerical tests confirm our analysis.
Keywords:
Incompressible Navier-Stokes equations with no-slip conditions Projection methods Energy stability Scalar auxiliary variable Fourth-order accuracy GePUP.
MSC:
Primary 76D05 65M20
journal:
1 Introduction
The incompressible Navier-Stokes equations (INSE) with no-slip conditions govern an enormous range of real-world phenomena such as blood flow, turbulence, atmosphere and ocean currents; they read
$\frac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot\nabla\mathbf{u}$ $=\mathbf{g}-\nabla p+\nu\Delta\mathbf{u}\quad\text{ in }\Omega,$ (1a)
$\nabla\cdot\mathbf{u}$ $=0\qquad\qquad\qquad\quad\text{ in }\Omega,$ (1b)
$\mathbf{u}$ $=\mathbf{0}\qquad\qquad\qquad\quad\text{ on }\partial\Omega,$ (1c)
where $t\in[0,+\infty)$ is time, $\Omega$ a domain, i.e., a bounded connected open subset of $\mathbb{R}^{{\scriptsize\textup{D}}}$, $\partial\Omega$ the domain boundary, $\mathbf{g}$ the external force, $p$ the pressure, $\mathbf{u}$ the velocity, and $\nu$ the kinematic viscosity.
In addition to efficiently, accurately, and faithfully reproducing the physical processes modeled by the equations, we confront four major challenges (FMC) of numerically solving the INSE,
  1. (FMC-1)
    How to fulfill the solenoidal condition (1b) and other physical constraints such as the monotonic decrease of the kinetic energy?
  2. (FMC-2)
    How to ensure various types of numerical stability?
  3. (FMC-3)
    How to obtain high-order convergence both in time and in space?
  4. (FMC-4)
    How to decouple time integration from spatial discretization so that (i) the entire solver is constituted by orthogonal modules for these aspects, and (ii) solution methods for each aspect can be employed in a black-box manner and thus easily changed to make the entire INSE solver versatile?
(FMC-1) concerns the prominent feature of mass conservation: neither source nor sink exists anywhere inside the domain. It is well known that a violation of (1b), even with small errors, might lead to qualitatively different flow patterns, especially for large Reynolds numbers. Another important physical constraint to be fulfilled in this work is the monotonic decrease of the total kinetic energy as defined in (44).
In (FMC-2), a crucial and indispensable type of numerical stability is the eigenvalue stability for the main evolutionary variable, which is typically the velocity. In addition, preserving the monotonicity of kinetic energy is equivalent to ensuring numerical stability on the velocity with respect to the 2-norm.
Challenge (FMC-3) concerns accuracy and efficiency. Near no-slip boundaries, flows at high Reynolds numbers tend to develop structures of multiple length scales and time scales. A numerical method should resolve all scales that are relevant to the important physics. Compared with fourth- and higher-order methods, first- and second-order methods have simpler algorithms and cheaper computations, but towards a given accuracy the computational resources may be rapidly exhausted. It is shown both theoretically and numerically in (zhang16:_GePUP, , Sec. 7) that fourth-order methods may have a large efficiency advantage over second-order methods.
Challenge (FMC-3) also concerns faithfully simulating flows where velocity derivatives such as vorticity crucially affect the physics. For first-order finite volume/difference methods, the computed velocity converges, but the vorticity does not, nor does the velocity gradient tensor. Consequently, the $O(1)$ error in $\nabla\mathbf{u}$ may lead to structures different from that of the original flow. In other words, it is not clear whether or not solutions of a first-order method have converged to the right physics. Similar suspicions apply to second-order methods for flows where second derivatives of the velocity are important.
Challenge (FMC-4) concerns versatility and user-friendliness of the numerical solver. To cater for the problem at hand, it is often desirable to change the time integrator from one to another. For example, flows with high viscosity are usually stiff while those with small viscosity are not; accordingly, an implicit time integrator should be used in the former case while an explicit one is usually suitable for the latter. If the internal details of a time integrator are coupled into the INSE solver in a boilerplate manner, it would be difficult and very inconvenient to change the time integrator; see also the discussion in the paragraph under (4). Hence a time integrator should be treated as a black box: for the ordinary differential equation (ODE) $\frac{\mathrm{d}U}{\mathrm{d}t}=f(U,t)$, we should only need to feed into the time integrator the initial condition $U^{n}$ and samples of $f$ at a number of time instances to get the solution $U^{n+1}$ from the black box.
This versatility further leads to user-friendliness. Analogous to orthogonal bases of a vector space, the mutually independent policies span a space of solvers, where each solver can be conveniently assembled by selecting a module for each constituting policy. For example, a specific INSE solver is formed by choosing semi-implicit Runge-Kutta (RK) for time integration, finite volume for spatial discretization, fourth-order for accuracy and so on; see Table 1.
1.1 Previous methods related to this work
In the original projection method independently proposed by Chorin Chorin_1968_numerical and Temam temam69:_sur_navier_stokes , the initial condition $\mathbf{u}^{n}\approx\mathbf{u}(t^{n})$ is first advanced to an auxiliary velocity $\mathbf{u}^{*}$ without worrying about the pressure gradient term and then $\mathbf{u}^{*}$ is projected to the divergence-free space to obtain $\mathbf{u}^{n+1}$,
$\frac{\mathbf{u}^{*}-\mathbf{u}^{n}}{k}$ $=-\mathbf{C}(\mathbf{u}^{*},\mathbf{u}^{n})+\mathbf{g}^{n}+\nu{\mathbf{L}}\mathbf{u}^{*},$ (2a)
$\mathbf{u}^{n+1}$ $={\mathbf{P}}\mathbf{u}^{*},$ (2b)
where $k$ is the time step size, $t^{n}$ the starting time of the $n$th step, $\mathbf{g}^{n}\approx\mathbf{g}(t^{n})$, $\mathbf{C}(\mathbf{u}^{*},\mathbf{u}^{n})\approx[(\mathbf{u}\cdot\nabla)\mathbf{u}](t^{n})$, and $\mathbf{L}$} and $\mathbf{P}$} discrete approximations of the Laplacian $\Delta$ and the Leray-Helmholtz projection $\mathscr{P}$}, respectively; see Section 2.1.
1.1.1 Second-order methods with fractional time stepping
The original projection method is first-order accurate and its improvement to the second order has been the aim of many subsequent works; see, e.g., Kim.Moin_1985_application ; bell89:_secon_order_projec_method_incom ; orszag86:_bound_condit_incom_flows ; e.03:_gauge ; guermond06:_overview_of_projection_methods_for_incompressible_flows ; brown01:_accurate_projection_methods_for_ins and references therein. A common basis of many second-order methods is the temporal discretization of (1) with the trapezoidal rule,
$\frac{\mathbf{u}^{n+1}-\mathbf{u}^{n}}{k}+\nabla p^{n+\frac{1}{2}}$ $=-[(\mathbf{u}\cdot\nabla)\mathbf{u}]^{n+\frac{1}{2}}+\mathbf{g}^{n+\frac{1}{2}}+\frac{\nu}{2}\Delta(\mathbf{u}^{n+1}+\mathbf{u}^{n}),$ (3a)
$\nabla\cdot\mathbf{u}^{n+1}$ $=0,$ (3b)
where $p^{n+\frac{1}{2}}\approx p(t^{n+\frac{1}{2}})$, $\mathbf{g}^{n+\frac{1}{2}}\approx\mathbf{g}(t^{n+\frac{1}{2}})$, and $(\mathbf{u}\cdot\nabla)\mathbf{u}]^{n+\frac{1}{2}}\approx[(\mathbf{u}\cdot\nabla)\mathbf{u}](t^{n+\frac{1}{2}})$ are numerical approximations at $t^{n+\frac{1}{2}}:=\frac{1}{2}(t^{n}+t^{n+1})$.
Replacing the gradient $\nabla$, the divergence $\nabla\cdot$, and the Laplacian $\Delta$ in (3) respectively with their second-order discrete counterparts $\mathbf{G}$}, $\mathbf{D}$}, and $\mathbf{L}$} yields
$A\begin{bmatrix}\mathbf{u}^{n+1}\\ p^{n+\frac{1}{2}}\end{bmatrix}:=\begin{bmatrix}\frac{1}{k}{\mathbf{I}}-\frac{\nu}{2}{\mathbf{L}}&{\mathbf{G}}\\ -{\mathbf{D}}&\mathbf{0}\end{bmatrix}\begin{bmatrix}\mathbf{u}^{n+1}\\ p^{n+\frac{1}{2}}\end{bmatrix}=\mathbf{F}.$ (4)
Since ${\mathbf{G}}^{T}=-{\mathbf{D}}$, the matrix $A$ has a saddle point structure and the above method is often called the saddle point approach. Despite its simplicity, this approach has two main disadvantages. First, the spatial discretization and time integration are coupled in a boilerplate manner and thus a change of either part would demand a complete rederivation of the matrix $A$. For fourth- or higher-order accuracy in time integration, it is often too complicated to have an explicit expression of the matrix $A$, as $A$ contains all internal details of the time integrator. Consequently, it is highly difficult for this approach to address challenges (FMC-3,4). Second, it is challenging benzi05:_numer to efficiently solve the linear system (4) since all the velocity components and the pressure are coupled into a big unknown vector; in contrast, Chorin’s projection method only requires the solutions of linear systems with the unknowns as either the pressure or a velocity component.
In the fractional-stepping projection methods Kim.Moin_1985_application ; bell89:_secon_order_projec_method_incom ; brown01:_accurate_projection_methods_for_ins , one replaces $\mathbf{u}^{n+1}$ and $p^{n+\frac{1}{2}}$ in (3a) respectively with $\mathbf{u}^{*}$ and $q$, solves for the auxiliary velocity $\mathbf{u}^{*}$ with some boundary condition $\mathbf{B}(\mathbf{u}^{*})=\mathbf{0}$, obtains $\mathbf{u}^{n+1}$ by the projection $\mathbf{u}^{n+1}=\mathbf{u}^{*}-k\nabla\phi^{n+1}$, and updates the pressure with $p^{n+\frac{1}{2}}=q+\mathbf{U}(\phi^{n+1})$.
Fractional-stepping projection methods have been very successful. However, the choices of $q$, $\mathbf{B}(\mathbf{u}^{*})$, and $\mathbf{U}(\phi^{n+1})$ are coupled according to internal details of the time integrator brown01:_accurate_projection_methods_for_ins . Consequently, switching from one time integrator to another calls for a new derivation. Furthermore, although appearing divorced, the velocity and the pressure are still implicitly coupled by the boundary condition of $\mathbf{u}^{*}$, with the coupling determined not by physics but still by internal details of the time integrator. Hence these methods are not suitable for tackling the challenges (FMC-3,4) either.
1.1.2 The formulation via the pressure Poisson equation (PPE)
As a specialization of Newton’s second law, the momentum equation (1a) can be rewritten as
$\mathbf{a}^{*}=\mathbf{a}+\nabla p,$ (5)
where the Eulerian accelerations are vector functions
$\mathbf{a}:=\frac{\partial\mathbf{u}}{\partial t},\qquad\mathbf{a}^{*}:=-\mathbf{u}\cdot\nabla\mathbf{u}+\mathbf{g}+\nu\Delta\mathbf{u}.$ (6)
The PPE describes an instantaneous relation between the pressure and the velocity in the INSE and, on no-slip domains, has the form
$\Delta p$ $=\nabla\cdot(\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u}+\nu\Delta\mathbf{u})\qquad\text{ in }\Omega,$ (7a)
$\mathbf{n}\cdot\nabla p$ $=\mathbf{n}\cdot(\mathbf{g}+\nu\Delta\mathbf{u})\quad\qquad\qquad\quad\text{ on }\partial\Omega,$ (7b)
where (7b) follows from the normal component of (1a) and the no-slip conditions (1c) while (7a) from the divergence of (1a) and the divergence-free condition (1b). For the PPE with other boundary conditions, (7b) should be replaced with the normal component of (5). As explained in Section 2.1, the pressure gradient is uniquely determined from $\mathbf{a}^{*}$ by $\nabla p=({\mathcal{I}}-{\mathscr{P}})\mathbf{a}^{*}$ where $\mathcal{I}$} and $\mathscr{P}$} are the identity operator and the Leray-Helmholtz projection, respectively. Thus neither the initial condition nor the boundary condition of the pressure $p$ is needed in the INSE.
(1a), (1c), (7), and the additional boundary condition $\nabla\cdot\mathbf{u}=0$ on $\partial\Omega$ are collectively called the PPE formulation of the INSE on no-slip domains gresho87:_navier_stokes . In terms of computation, however, the PPE formulation has a decisive advantage over the original INSE. If (1a) is discretized in time with (1b) as a constraint, the resulting index-2 differential algebraic system may suffer from large order reductions sanderse12:_accur_runge_kutta_navier_stokes . In contrast, replacing the divergence-free constraint with the PPE avoids this difficulty. As such, the PPE formulation allows the time integrator to be treated as a black box and thus to be easily changed; indeed, the pressure is an implicit function of $\mathbf{u}$ and its interaction with $\mathbf{u}$ is completely decoupled from internal details of the time integrator. Also, there is no need to introduce nonphysical auxiliary variables. These advantages of the PPE formulation lead to quite a number of successful numerical methods kleiser80:_treat_incom_bound_condit_d ; gresho87:_navier_stokes ; henshaw94:_fourt_accur_method_incom_navier ; johnston04:_accur_navier_stokes ; liu07:_stabil_conver_effic_navier_stokes ; liu10_stable_accurate_pressure_unsteady_incompressible_viscous_flow ; shirokoff11:_navier_stokes .
Unfortunately, as observed by Liu, Liu & Pego liu07:_stabil_conver_effic_navier_stokes , (1a) and (7a) yield
$\frac{\partial\nabla\cdot\mathbf{u}}{\partial t}=0;$ (8)
this degenerate equation implies that in the PPE formulation we have no control over $\nabla\cdot\mathbf{u}$ and its evolution is up to the particularities of the numerical schemes. Our tests show that a fourth-order finite-volume method-of-lines (MOL) discretization of the PPE formulation is unstable, with the computed velocity divergence growing indefinitely near the domain boundary.
1.1.3 The formulation of unconstrained PPE (UPPE)
The application of the Leray-Helmholtz projection $\mathscr{P}$} to (5) yields
$\frac{\partial\mathbf{u}}{\partial t}-{\mathscr{P}}\mathbf{a}^{*}=\nu\nabla(\nabla\cdot\mathbf{u}),$ (9)
where the zero right-hand side (RHS) is added for stability reasons liu07:_stabil_conver_effic_navier_stokes . The divergence of (9) and the second identity in (20) give
$\frac{\partial(\nabla\cdot\mathbf{u})}{\partial t}=\nu\Delta(\nabla\cdot\mathbf{u}),$ (10)
which, by the maximum principle of the heat equation, dictates an exponential decay of $\nabla\cdot\mathbf{u}$ in $\Omega$. A juxtaposition of (10) and (8) exposes a prominent advantage of (9) that any divergence residue is now well under control.
Via the identity $\nabla(\nabla\cdot\mathbf{u})=\Delta({\mathcal{I}}-{\mathscr{P}})\mathbf{u}$ and the Laplace-Leray commutator $\Delta,{\mathscr{P}}]:=\Delta{\mathscr{P}}-{\mathscr{P}}\Delta$, cf. Section 2.2, Liu, Liu & Pego liu07:_stabil_conver_effic_navier_stokes rewrote (9) as
$\frac{\partial\mathbf{u}}{\partial t}+{\mathscr{P}}(\mathbf{u}\cdot\nabla\mathbf{u}-\mathbf{g})+\nu[\Delta,{\mathscr{P}}]\mathbf{u}=\nu\Delta\mathbf{u},$ (11)
which provides a fresh viewpoint of the INSE as a controlled perturbation of the vector diffusion equation $\frac{\partial\mathbf{u}}{\partial t}=\nu\Delta\mathbf{u}$. For $\mathbf{u}\in H^{2}\cap H_{0}^{1}(\Omega,\mathbb{R}^{{\scriptsize\textup{D}}})$ with ${\mathcal{C}}^{3}$ boundary $\partial\Omega$, they gave a sharp bound on $\|[\Delta,{\mathscr{P}}]\mathbf{u}\|$ in terms of $\|\Delta\mathbf{u}\|$ and proved the unconditional stability and convergence of a first-order scheme,
$\left\langle\nabla p^{n},\nabla\phi\right\rangle$ $=\left\langle\mathbf{g}^{n}-\mathbf{u}^{n}\cdot\nabla\mathbf{u}^{n}+\nu\Delta\mathbf{u}^{n}-\nu\nabla\nabla\cdot\mathbf{u}^{n},\nabla\phi\right\rangle,$ (12a)
$\frac{\mathbf{u}^{n+1}-\mathbf{u}^{n}}{k}+\nabla p^{n}$ $=\mathbf{g}^{n}-\mathbf{u}^{n}\cdot\nabla\mathbf{u}^{n}+\nu\Delta\mathbf{u}^{n+1}\qquad\text{ in }\Omega,$ (12b)
$\mathbf{u}^{n+1}$ $=\mathbf{0}\qquad\qquad\qquad\qquad\qquad\qquad\text{ on }\partial\Omega,$ (12c)
where (12a) is the PPE in weak form with $\phi\in H^{1}(\Omega)$ and $\left\langle\mathbf{u},\mathbf{v}\right\rangle:=\int_{\Omega}\mathbf{u}\cdot\mathbf{v}\,\mathrm{d}V$.
Based on the UPPE formulation, a slip-corrected projection method liu10_stable_accurate_pressure_unsteady_incompressible_viscous_flow is developed with third-order accuracy both in time and in space.
From (11) and the contents in Sections 2.2 and 2.3, a strong form of UPPE can be deduced liu10_stable_accurate_pressure_unsteady_incompressible_viscous_flow as
$\frac{\partial\mathbf{u}}{\partial t}+\mathbf{u}\cdot\nabla\mathbf{u}$ $=\mathbf{g}-\nabla p+\nu\Delta\mathbf{u}\qquad\qquad\qquad\text{ in }\Omega,$ (13a)
$\mathbf{u}$ $=\mathbf{0}\qquad\qquad\qquad\qquad\qquad\qquad\text{ on }\partial\Omega,$ (13b)
$\Delta p$ $=\nabla\cdot(\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u})\qquad\qquad\quad\ \ \text{ in }\Omega,$ (13c)
$\mathbf{n}\cdot\nabla p$ $=\mathbf{n}\cdot(\mathbf{g}+\nu\Delta\mathbf{u}-\nu\nabla\nabla\cdot\mathbf{u})\qquad\text{ on }\partial\Omega.$ (13d)
The PPE (7) and the UPPE (13c,d) have slightly different forms and nonetheless a crucial distinction: it follows from the divergence of (13a) that (13c) leads to (10) whereas (7a) leads to (8).
Unfortunately, (13) is not yet suitable for the design of MOL-type finite volume and finite difference methods, due to two main reasons (TMR).
  1. (TMR-1)
    The Leray-Helmholtz projection is absent in (13) and thus any projection on the velocity in an MOL algorithm would be a mismatch of the numerical algorithm to the governing equations. Of course one can replace the velocity $\mathbf{u}$ with ${\mathscr{P}}\mathbf{u}$ in (13), but which should be replaced? In other words, which $\mathbf{u}$’s in (13) should be projected in MOL?
  2. (TMR-2)
    It is difficult for a discrete projection $\mathbf{P}$ with fourth- and higher-order accuracy to satisfy all properties of the Leray-Helmholtz projection $\mathscr{P}$} in (20). In particular, the discretely projected velocity may not be divergence-free. Then how does the approximation error of $\mathbf{P}$ to $\mathscr{P}$} affect the stability of the ODE system under the MOL framework? It is neither clear nor trivial how to answer this question with (13).
1.1.4 Generic projection and unconstrained PPE (GePUP)
A generic projection is a linear operator $\mathcal{P}$} on a vector space satisfying
${\mathcal{P}}\mathbf{u}=\mathbf{w}:=\mathbf{u}-\nabla\phi,$ (14)
where $\phi$ is a scalar function and $\nabla\cdot\mathbf{w}=0$ may or may not hold. Since $\phi$ is not specified in terms of $\mathbf{w}$, (14) is not a precise definition of $\mathcal{P}$}, but rather a characterization of a family of operators, which, in particular, includes the Leray-Helmholtz projection $\mathscr{P}$}. $\mathcal{P}$} can be used to perturb $\mathbf{u}$ to some non-solenoidal velocity $\mathbf{w}$ and is thus more flexible than $\mathscr{P}$} in characterizing discrete projections that fail to fulfill the divergence-free constraint exactly.
To accommodate the fact that the discrete velocity might not be divergence-free, we switch the evolutionary variable to a non-solenoidal velocity $\mathbf{w}={\mathcal{P}}\mathbf{u}$ instead of the divergence-free velocity $\mathbf{u}$ in the UPPE formulation (13). More precisely, the evolutionary variable $\mathbf{u}$ in the time-derivative term $\frac{\partial\mathbf{u}}{\partial t}$ is perturbed to a non-solenoidal velocity $\mathbf{w}:=\mathbf{u}-\nabla\phi$ where $\phi$ is some scalar function; meanwhile in the diffusion term we change $\mathbf{u}$ to $\mathbf{w}$ to set up a mechanism that drives the divergence towards zero. Then, there is no need to worry about the influence of $\nabla\cdot\mathbf{w}\neq 0$ on numerical stability because the evolution of $\mathbf{w}$ is not subject to the divergence-free constraint. These ideas lead to the GePUP formulation zhang16:_GePUP :
$\frac{\partial\mathbf{w}}{\partial t}$ $=\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u}-\nabla q+\nu\Delta\mathbf{w}$ $\text{in }\Omega,$ (15a)
$\mathbf{w}$ $=\mathbf{0},\ \ \mathbf{u}\cdot\bm{\tau}=0,\ \$$\text{on }\partial\Omega,$ (15b)
$\mathbf{u}$ $={\mathscr{P}}\mathbf{w}$ $\text{in }\Omega,$ (15c)
$\mathbf{u}\cdot\mathbf{n}$ $=0$ $\text{on }\partial\Omega,$ (15d)
$\Delta q$ $=\nabla\cdot(\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u})$ $\text{in }\Omega,$ (15e)
$\mathbf{n}\cdot\nabla q$ $=\mathbf{n}\cdot(\mathbf{g}+\nu\Delta\mathbf{u}-\nu\nabla\nabla\cdot\mathbf{u})$ $\text{on }\partial\Omega,$ (15f)
where (15e) and the divergence of (15) yield $\frac{\partial\left(\nabla\cdot\mathbf{w}\right)}{\partial t}=\nu\Delta\left(\nabla\cdot\mathbf{w}\right)$. Then either $\nabla\cdot\mathbf{w}=0$ or $\mathbf{n}\cdot\nabla\nabla\cdot\mathbf{w}=0$ on $\partial\Omega$ drives $\nabla\cdot\mathbf{w}$ towards zero.
1.2 The contribution of this work
We couple GePUP with electric boundary conditions shirokoff11:_navier_stokes ; rosales21:_high_poiss_navier_stokes and a scalar auxiliary variable (SAV) shen18:SAV ; shen19:SAV_Review to propose GePUP-E and GePUP-ES, variants of GePUP that enforce the solenoidal conditions, preserve energy stability, decouple time integration from spatial discretization, and lead to versatile algorithms that are fourth-order accurate both in time and in space. The letter ‘E’ in the acronyms stands for the electric boundary conditions while the letter ‘S’ for the SAV approach; altogether ‘ES’ also stands for energy stability.
GePUP-E and GePUP-ES answer all the challenges in (FMC-1,2,3,4).
  1. (A)
    We reformulate the INSE of two variables into the GePUP-E formulation (32) of three variables, embedding the solution manifold $\mathcal{M}(\mathbf{u},p)$ of the INSE in the higher-dimensional solution manifold $\mathcal{N}(\mathbf{w},\mathbf{u},q)$ of GePUP-E. More importantly, we equip this embedding with a divergence-decaying mechanism that drives any deviating solution in $\mathcal{N}(\mathbf{w},\mathbf{u},q)$ back to $\mathcal{M}(\mathbf{u},p)$. We prove the equivalence of INSE and GePUP-E, the convergence of the non-solenoidal velocity $\mathbf{w}$ to the divergence-free velocity $\mathbf{u}$, the exponential decay of the divergence $\nabla\cdot\mathbf{w}$, and the monotonic decrease of the kinetic energy. As such, GePUP-E resolves the difficulties in (TMR-1,2).
  2. (B)
    By coupling GePUP-E to SAV shen18:SAV ; shen19:SAV_Review , we propose the GePUP-ES formulation in (47), design a family of semi-discrete GePUP-ES algorithms, and prove their energy stability in Theorem 5.4.
  3. (C)
    Based on (B), we further propose a family of fully discrete INSE solvers, named GePUP-ES-SDIRK, to answer all challenges in (FMC-1,2,3,4).
(A), (B), and (C) are elaborated in Sections 3, 4, and 5, respectively. In Section 2, we introduce notation to make this paper somewhat self-contained. We test GePUP-ES-SDIRK in Section 6 and draw conclusions in Section 7.
2 Preliminaries
Throughout this paper, we denote by $\left\langle\cdot,\cdot\right\rangle$ the $L^{2}$ inner product of vector- (or scalar-) valued functions $\mathbf{u}$ and $\mathbf{v}$ over $\Omega$, $\left\langle\mathbf{u},\mathbf{v}\right\rangle:=\int_{\Omega}\mathbf{u}\cdot\mathbf{v}\,\mathrm{d}V$, and by $\|\cdot\|$ the induced $L^{2}$ norm $\|\mathbf{u}\|:=\sqrt{\left\langle\mathbf{u},\mathbf{u}\right\rangle}$.
We start with a well-known result on boundary value problems (BVPs).
Theorem 2.1 (Solvability of BVPs with pure Neumann conditions)
Suppose $f$ and $g$ are two sufficiently smooth functions. Then there exists a unique solution (up to an additive constant) for the Neumann BVP
$\Delta\phi$ $=f\qquad\text{ in }\Omega;$ (16a)
$\mathbf{n}\cdot\nabla\phi$ $=g\qquad\text{ on }\partial\Omega$ (16b)
if and only if $\int_{\Omega}f\,\mathrm{d}V=\int_{\partial\Omega}g\,\mathrm{d}A$.
Proof
See (taylor11:_partial_differ_equat_i, , page 409). ∎
2.1 The Leray-Helmholtz projection $\mathscr{P}$}
Theorem 2.2 (Helmholtz decomposition)
A continuously differentiable vector field $\mathbf{v}^{*}$ in a domain $\Omega$ can be uniquely decomposed into a divergence-free part $\mathbf{v}$ and a curl-free part $\nabla\phi$,
$\mathbf{v}^{*}=\mathbf{v}+\nabla\phi,$ (17)
where $\mathbf{v}\cdot\mathbf{n}$ is given a priori on $\partial\Omega$ and satisfies $\oint_{\partial\Omega}\mathbf{v}\cdot\mathbf{n}=0$.
Proof
The decomposition can be realized by solving
$\left\{\begin{array}[]{rll}\Delta\phi&=\nabla\cdot\mathbf{v}^{*}&\text{ in }\Omega,\\ \mathbf{n}\cdot\nabla\phi&=\mathbf{n}\cdot(\mathbf{v}^{*}-\mathbf{v})&\text{ on }\partial\Omega,\end{array}\right$. (18)
since Theorem 2.1 uniquely determines $\nabla\phi$ with $\oint_{\partial\Omega}\mathbf{v}\cdot\mathbf{n}=0$. ∎
The Leray-Helmholtz projection $\mathscr{P}$} is an idempotent operator that maps a vector field $\mathbf{v}^{*}$ to its divergence-free part $\mathbf{v}$, c.f. the decomposition (17), i.e.,
${\mathscr{P}}\mathbf{v}^{*}:=\mathbf{v}=\mathbf{v}^{*}-\nabla\phi.$ (19)
The proof of Theorem 2.2 implies the constructive form ${\mathscr{P}}={\mathcal{I}}-\nabla(\Delta_{n})^{-1}\nabla\cdot$, where $(\Delta_{n})^{-1}$ denotes solving (18). For a ${\mathcal{C}}^{1}$ vector field $\mathbf{v}^{*}$ and a ${\mathcal{C}}^{1}$ scalar field $\phi$, we have
${\mathscr{P}}^{2}={\mathscr{P}},\quad\nabla\cdot{\mathscr{P}}\mathbf{v}^{*}=0,\quad{\mathscr{P}}\nabla\phi=\mathbf{0}.$ (20)
2.2 The Laplace-Leray commutator $\Delta{\mathscr{P}}-{\mathscr{P}}\Delta$
On periodic domains, $\Delta$ and $\mathscr{P}$} commute. However, one main difficulty for no-slip domains is the fact that $\Delta{\mathscr{P}}-{\mathscr{P}}\Delta\neq\mathbf{0}$. In this subsection we rephrase several results in liu07:_stabil_conver_effic_navier_stokes .
Lemma 2.3
The divergence-gradient commutator defined as
${\mathcal{B}}=[\nabla\cdot,\nabla]:=\Delta$-$\nabla\nabla\cdot$ (21)
satisfies $\nabla\cdot{\mathcal{B}}=0$, $\Delta{\mathscr{P}}={\mathcal{B}}$, and in three dimensions $\mathcal{B}$$\nabla\times\nabla\times$.
Proof
$\nabla\cdot{\mathcal{B}}=0$ follows from (21) and $\Delta\nabla\cdot=\nabla\cdot\Delta$ while ${\mathcal{B}}\mathbf{v}^{*}=-\nabla\times\nabla\times\mathbf{v}^{*}$ from the tensor notation and the epsilon-delta relation. $\Delta{\mathscr{P}}={\mathcal{B}}$ holds because
$\Delta{\mathscr{P}}\mathbf{v}^{*}=\Delta(\mathbf{v}^{*}-\nabla\phi)=\Delta\mathbf{v}^{*}-\nabla\Delta\phi=\Delta\mathbf{v}^{*}-\nabla\nabla\cdot\mathbf{v}^{*},$ (22)
where we have applied (19), the commutativity of $\Delta$ and $\nabla$, and (21). ∎
With ${\mathscr{P}}\nabla\phi=\mathbf{0}$ in (20), the first and third terms in (22) lead to
$\mathscr{P}\Delta\mathscr{P}=\mathscr{P}\Delta$ (23)
because $\mathscr{P}\Delta\mathscr{P}\mathbf{v}^{*}=\mathscr{P}\Delta\mathbf{v}^{*}$ holds for any sufficiently smooth vector field $\mathbf{v}^{*}$.
Then Lemma 2.3 and (23) give
Corollary 2.4
The Laplace-Leray commutator is
$\Delta,{\mathscr{P}}]:=\Delta{\mathscr{P}}-{\mathscr{P}}\Delta=({\mathcal{I}}-{\mathscr{P}})\Delta{\mathscr{P}}=({\mathcal{I}}-{\mathscr{P}}){\mathcal{B}}=-({\mathcal{I}}-{\mathscr{P}})(\nabla\times\nabla\times),$ (24)
where $\mathcal{I}$} is the identity operator and the last equality holds only in three dimensions.
2.3 The Stokes pressure
By (24), the action of the Laplace-Leray commutator on any vector field $\mathbf{v}^{*}$ results in the gradient of some scalar field. In the case of $\mathbf{v}^{*}$ being the velocity $\mathbf{u}$ in the INSE, the scalar is known as the Stokes pressure liu07:_stabil_conver_effic_navier_stokes :
$\nabla p_{s}:=(\Delta{\mathscr{P}}-{\mathscr{P}}\Delta)\mathbf{u}.$ (25)
It follows from (24) and (25) that
${\mathcal{B}}\mathbf{u}={\mathscr{P}}{\mathcal{B}}\mathbf{u}+\nabla p_{s}.$
Then (20) and $\nabla\cdot{\mathcal{B}}=0$ in Lemma 2.3 yield $\Delta p_{s}=0$, i.e., the Stokes pressure is harmonic. Interestingly, the vector field $\nabla p_{s}$ is both divergence-free and curl-free.
Define another scalar $p_{c}$ as
$\nabla p_{c}:=({\mathcal{I}}-{\mathscr{P}})\left(\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u}\right).$ (26)
Apply the Leray-Helmholtz projection to (1a), use the commutator (24), invoke the definitions (25) and (26), and we have
$\nabla p=\nabla p_{c}+\nu\nabla p_{s}.$
The pressure gradient in the INSE consists of two parts: $\nabla p_{c}$ balances the divergence of the forcing term and the nonlinear convection term while $\nabla p_{s}$ accounts for the Laplace-Leray commutator. In the two limiting cases of $\nu\rightarrow 0$ and $\nu\rightarrow+\infty$, the pressure gradient is dominated by $\nabla p_{c}$ and $\nu\nabla p_{s}$, respectively.
2.4 Vector identities
See a standard text on differentiable manifolds such as munkres97:_analy_manif for a proof of
Theorem 2.5 (Gauss-Green)
A scalar or vector function $u\in\mathcal{C}^{1}(\overline{\Omega})$ satisfies
$\int_{\Omega}\frac{\partial u}{\partial x_{i}}\mathrm{d}V=\int_{\partial\Omega}un_{i}\,\mathrm{d}A,$
where $n_{i}$ is the $i$th component of the unit outward normal $\mathbf{n}$ of $\partial\Omega$.
Apply Theorem 2.5 to $uv$ and we have
Lemma 2.6 (Integration-by-parts)
For $u,v\in\mathcal{C}^{1}(\overline{\Omega})$, we have
$\int_{\Omega}\frac{\partial u}{\partial x_{i}}v\,\mathrm{d}V=-\int_{\Omega}u\frac{\partial v}{\partial x_{i}}\mathrm{d}V+\int_{\partial\Omega}uvn_{i}\,\mathrm{d}A,$ (27)
where $n_{i}$ is the $i$th component of the unit outward normal $\mathbf{n}$ of $\partial\Omega$.
Replace $v$ in (27) with $\frac{\partial v}{\partial x_{i}}$, sum over $i$, and we have
Lemma 2.7 (Green’s formula)
For $u,v\in\mathcal{C}^{2}(\overline{\Omega})$, we have
$\int_{\Omega}\nabla u\cdot\nabla v\,\mathrm{d}V=-\int_{\Omega}u\Delta v\,\mathrm{d}V+\int_{\partial\Omega}u\frac{\partial v}{\partial\mathbf{n}}\mathrm{d}A.$ (28)
2.5 B-stable and algebraically stable RK methods
To solve an ODE system $\mathbf{u}^{\prime}=\mathbf{f}(\mathbf{u},t)$, an $s$-stage RK method is a one-step method of the form
$\left\{\begin{array}[]{l}\mathbf{y}_{i}=\mathbf{f}(\mathbf{U}^{n}+k\sum_{j=1}^{s}a_{i,j}\mathbf{y}_{j},t_{n}+c_{i}k),\\ \mathbf{U}^{n+1}=\mathbf{U}^{n}+k\sum_{j=1}^{s}b_{j}\mathbf{y}_{j},\end{array}\right$. (29)
where $i=1,2,\ldots,s$ and the coefficients $a_{i,j}$, $b_{j}$, $c_{i}$ are real.
A function $\mathbf{f}:\mathbb{R}^{{\scriptsize\textup{D}}}\times[0,+\infty)\to\mathbb{R}^{{\scriptsize\textup{D}}}$ is one-sided Lipschitz continuous if
$\forall t\geq 0,\forall\mathbf{u},\mathbf{v}\in\mathbb{R}^{{\scriptsize\textup{D}}},\ \ \left\langle\mathbf{u}-\mathbf{v},\ \mathbf{f}(\mathbf{u},t)-\mathbf{f}(\mathbf{v},t)\right\rangle\leq\mu\|\mathbf{u}-\mathbf{v}\|^{2},$ (30)
where $\mu$ is the one-sided Lipschitz constant of $\mathbf{f}$. The ODE system is contractive or monotone if $\mathbf{f}$ satisfies (30) with $\mu=0$.
A contractive ODE system is dissipative: for any solutions $\mathbf{u}(t)$ and $\mathbf{v}(t)$, the norm $\|\mathbf{u}(t)-\mathbf{v}(t)\|$ decreases monotonically as $t$ increases. In other words, different solution trajectories of a contractive ODE system never depart from each other, and hence small perturbations remain small. This leads to
Definition 2.8 (B-stability butcher75:_runge_kutta )
A one-step method is B-stable if, for any contractive ODE system, each pair of numerical solutions $\mathbf{U}^{n}$ and $\mathbf{V}^{n}$ satisfy
$\forall n=0,1,\ldots,\quad\|\mathbf{U}^{n+1}-\mathbf{V}^{n+1}\|\leq\|\mathbf{U}^{n}-\mathbf{V}^{n}\|.$
It can be shown that B-stable methods are A-stable.
Definition 2.9
An RK method is algebraically stable if
  • the RK weights $b_{1},b_{2},\ldots,b_{s}$ are nonnegative,
  • the following symmetric matrix $M\in\mathbb{R}^{s\times s}$ is positive semidefinite:
    $m_{i,j}=b_{i}a_{i,j}+b_{j}a_{j,i}-b_{i}b_{j}.$ (31)
An algebraically stable RK method is B-stable and thus A-stable hairer96:_solvin_ordin_differ_equat_ii .
3 The GePUP-E formulation
Boundary conditions of the evolutionary variable $\mathbf{w}$ play a crucial role in establishing a stable numerical scheme. As mentioned in Subsection 1.1.4, the GePUP formulation (15) leads to the heat equation $\frac{\partial\left(\nabla\cdot\mathbf{w}\right)}{\partial t}=\nu\Delta\left(\nabla\cdot\mathbf{w}\right)$. However, neither the homogeneous Dirichlet nor the homogeneous Neumann condition is explicitly imposed on $\nabla\cdot\mathbf{w}$ in (15), thus it is difficult to prove the dacay of $\nabla\cdot\mathbf{w}$. To fix this glitch, we draw inspiration from the excellent work of Rosales, Shirokoff, and their colleagues shirokoff11:_navier_stokes ; rosales21:_high_poiss_navier_stokes to adapt “electric” boundary conditions into GePUP, proposing
Definition 3.1
The GePUP-E formulation of INSE on no-slip domains is
$\frac{\partial\mathbf{w}}{\partial t}$ $=\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u}-\nabla q+\nu\Delta\mathbf{w}$ $\text{in }\Omega,$ (32a)
$\mathbf{w}\cdot\bm{\tau}$ $=0,\ \ \nabla\cdot\mathbf{w}=0\ \$$\text{on }\partial\Omega,$ (32b)
$\mathbf{u}$ $={\mathscr{P}}\mathbf{w}$ $\text{in }\Omega,$ (32c)
$\mathbf{u}\cdot\mathbf{n}$ $=0$ $\text{on }\partial\Omega,$ (32d)
$\Delta q$ $=\nabla\cdot(\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u})$ $\text{in }\Omega,$ (32e)
$\mathbf{n}\cdot\nabla q$ $=\mathbf{n}\cdot(\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u}+\nu\Delta\mathbf{w})+\lambda\mathbf{n}\cdot\mathbf{w}$ $\text{on }\partial\Omega,$ (32f)
where $\mathbf{u}$ is the divergence-free velocity in (1), $\mathbf{w}=\mathbf{u}-\nabla\phi$ a non-solenoidal velocity for some scalar function $\phi$, $\mathbf{n}$ and $\bm{\tau}$ the unit normal and unit tangent vector of $\partial\Omega$, respectively, and $\lambda$ a nonnegative penalty parameter. The two velocities $\mathbf{w}$ and $\mathbf{u}$ have the same initial condition in $\overline{\Omega}$, the closure of $\Omega$, i.e.,
$\forall\mathbf{x}\in\overline{\Omega},\quad\mathbf{w}(\mathbf{x},t_{0})=\mathbf{u}(\mathbf{x},t_{0}).$ (33)
(32b) is different from (15b). First, the boundary condition $\mathbf{n}\cdot\mathbf{w}=0$ in (15b) is removed and the term $\lambda\mathbf{n}\cdot\mathbf{w}$ is added in (32f) so that any nonzero $\mathbf{n}\cdot\mathbf{w}$ decays exponentially towards zero; see Lemma 3.3. Second, the boundary condition $\nabla\cdot\mathbf{w}=0$ is added in (32b) to set up an exponential decay of $\nabla\cdot\mathbf{w}$; see Theorem 3.4. Third, the boundary condition $\mathbf{w}\cdot\bm{\tau}=0$ is added in (32b) to close the vector diffusion equation governing the evolution of $\mathbf{w}$. Lastly, the boundary condition $\mathbf{u}\cdot\bm{\tau}=0$ in (15b) is removed because, as a perturbed version of $\mathbf{u}$, the non-solenoidal velocity $\mathbf{w}$ is $\mathbf{u}$ when the initial condition (33) is imposed; see Lemma 3.5. Another initial condition weaker than (33) is
$\forall\mathbf{x}\in\partial{\Omega},\quad\mathbf{w}(\mathbf{x},t_{0})=\mathbf{u}(\mathbf{x},t_{0}).$ (34)
To connect $\mathbf{w}$ to $\mathbf{u}$, (32c) appears to be the most natural choice.
Compared to the formulation in shirokoff11:_navier_stokes ; rosales21:_high_poiss_navier_stokes , the GePUP-E formulation (32) facilitates the design and analysis of numerical schemes that treat the nonlinear convection term and the pressure gradient term explicitly; see Definition 5.1.
Lemma 3.2
The Neumann BVP (32c,d) admits a unique solution of $\mathbf{u}$.
Proof
(32c,d) and the definition $\mathbf{w}=\mathbf{u}-\nabla\phi$ yield
$\begin{array}[]{rl}\Delta\phi&=-\nabla\cdot\mathbf{w}\qquad\text{ in }\Omega,\\ \mathbf{n}\cdot\nabla\phi&=-\mathbf{n}\cdot\mathbf{w}\qquad\text{ on }\partial\Omega.\end{array}$ (35)
Then the proof is completed by Theorem 2.1 and the divergence theorem. ∎
3.1 The exponential decay of $\mathbf{n}\cdot\mathbf{w}$ and $\nabla\cdot\mathbf{w}$ in (32)
Although the no-penetration condition $\mathbf{n}\cdot\mathbf{w}=0$ is not explicitly stated in (32), the exponential decay of $\mathbf{n}\cdot\mathbf{w}$ on $\partial\Omega$ is guaranteed by
Lemma 3.3
The GePUP-E formulation (32) satisfies
$\frac{\partial(\mathbf{n}\cdot\mathbf{w})}{\partial t}=-\lambda\mathbf{n}\cdot\mathbf{w}\quad\mathnormal{on~}\partial\Omega$ (36)
and thus $\mathbf{n}\cdot\mathbf{w}(t)=e^{-\lambda(t-t_{0})}\mathbf{n}\cdot\mathbf{w}(t_{0})$ holds on $\partial\Omega$. In particular, we have
$\mathbf{n}\cdot\mathbf{w}(t_{0})=0\ \Rightarrow\ \forall t>t_{0},\ \mathbf{n}\cdot\mathbf{w}(t)=0.$
Proof
(36) follows immediately from (32f) and the normal component of (32a). ∎
The exponential decay of $\nabla\cdot\mathbf{w}$ is guaranteed by
Theorem 3.4
The GePUP-E formulation (32) satisfies
$\frac{\partial\left(\nabla\cdot\mathbf{w}\right)}{\partial t}=\nu\Delta\left(\nabla\cdot\mathbf{w}\right)\quad\text{ in }\Omega,$ (37)
which implies
$\frac{\mathrm{d}}{\mathrm{d}t}\|\nabla\cdot\mathbf{w}\|^{2}=-2\nu\left\|\nabla\left(\nabla\cdot\mathbf{w}\right)\right\|^{2},$ (38)
which further implies
$\|\nabla\cdot\mathbf{w}(t)\|\leq e^{-\nu C(t-t_{0})}\|\nabla\cdot\mathbf{w}(t_{0})\|$ (39)
where $t_{0}$ is the initial time and $C$ a positive constant independent of $\mathbf{w}$.
Proof
(37) follows from the divergence of (32a), (32e), and the commutativity of $\nabla\cdot$ and $\Delta$. Multiply (37) with $\nabla\cdot\mathbf{w}$, integrate over $\Omega$, and we have
$\begin{array}[]{rl}&\frac{1}{2}\frac{\mathrm{d}}{\mathrm{d}t}\|\nabla\cdot\mathbf{w}\|^{2}=\nu\left\langle\Delta\left(\nabla\cdot\mathbf{w}\right),\nabla\cdot\mathbf{w}\right\rangle\\ =&-\nu\left\|\nabla\left(\nabla\cdot\mathbf{w}\right)\right\|^{2}+\nu\oint_{\partial\Omega}\left(\nabla\cdot\mathbf{w}\right)\frac{\partial\left(\nabla\cdot\mathbf{w}\right)}{\partial\mathbf{n}}\mathrm{d}A=-\nu\left\|\nabla\left(\nabla\cdot\mathbf{w}\right)\right\|^{2},\end{array}$
where the second step follows from Green’s formula (28) with $u=v=\nabla\cdot\mathbf{w}$ and the last from (32b). Then the Poincaré-Friedrichs inequality and (38) imply $\frac{\mathrm{d}}{\mathrm{d}t}\|\nabla\cdot\mathbf{w}\|^{2}\leq-2\nu C\|\nabla\cdot\mathbf{w}\|^{2}$ and then (39) follows from
$\begin{array}[]{l}\frac{\mathrm{d}}{\mathrm{d}t}\left(e^{2\nu Ct}\|\nabla\cdot\mathbf{w}\|^{2}\right)=e^{2\nu Ct}\left(2\nu C\|\nabla\cdot\mathbf{w}\|^{2}+\frac{\mathrm{d}}{\mathrm{d}t}\|\nabla\cdot\mathbf{w}\|^{2}\right)\leq 0.\end{array}\quad\qed$
Due to the exponential decay in (39), the convergence of $\mathbf{w}$ to $\mathbf{u}$ should be sufficiently fast for most practical applications. In numerical simulations via GePUP-E, it is suggested to set the initial condition of $\mathbf{w}$ to that of $\mathbf{u}$. Then the mechanism of exponential decay in Theorem 3.4 will suppress divergence residue caused by truncation errors of spatial operators; see Theorem 3.7.
3.2 Recovering INSE from GePUP-E
Lemma 3.5
With the initial condition (33), the GePUP-E formulation (32) recovers the INSE in (1) for all $t\geq t_{0}$.
Proof
(33) and (39) yield $\nabla\cdot\mathbf{w}(t)=0$. With $\mathbf{w}\cdot\mathbf{n}=0$ in Lemma 3.3, the BVP (35) reduces to a Laplace equation with homogeneous Neumann conditions, for which $\phi$ being a constant in $\Omega$ is a particular solution. Thus $\nabla\phi=\mathbf{0}$ and $\mathbf{w}=\mathbf{u}$. ∎
Lemma 3.5 implies the well-posedness of (32e,f). In comparison to (15f), (32f) contains the convection term so that the well-posedness of the Neumann BVP (32e,f) is independent of the boundary condition of $\mathbf{u}$.
Lemma 3.6
With the initial condition (33), the Neumann BVP (32e,f) admits a unique solution of $\nabla q$.
Proof
The divergence theorem and the commutativity of $\nabla\cdot$ and $\Delta$ imply
$\oint_{\partial\Omega}\left(\mathbf{n}\cdot\nu\Delta\mathbf{w}+\lambda\mathbf{n}\cdot\mathbf{w}\right)\mathrm{d}A=\int_{\Omega}\left(\nu\Delta\nabla\cdot\mathbf{w}+\lambda\nabla\cdot\mathbf{w}\right)\mathrm{d}V=0,$
where the second step follows from Lemma 3.5. The divergence theorem gives
$\int_{\Omega}\nabla\cdot\left(\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u}\right)\mathrm{d}V=\oint_{\partial\Omega}\mathbf{n}\cdot(\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u})\,\mathrm{d}A.$
Then the rest of the proof follows from Theorem 2.1. ∎
Theorem 3.7
GePUP-E (32) with the initial condition (34) satisfies
$\forall\epsilon>0,\ \exists t^{*}>t_{0},\ \text{ s.t. }\forall t>t^{*},\quad\left\{\begin{array}[]{l}\sup_{\mathbf{x}\in\Omega}\left|\nabla\cdot\mathbf{w}(\mathbf{x},t)\right|<\epsilon\\ \left\|\mathbf{w}(\mathbf{x},t)-\mathbf{u}(\mathbf{x},t)\right\|<\epsilon,\\ \left\|q(\mathbf{x},t)-p(\mathbf{x},t)\right\|<\epsilon.\end{array}\right$.
Proof
The first limit follows from (39). The second limit then follows from arguments similar to those in the proof of Lemma 3.5. The third limit is a consequence of the second limit. ∎
3.3 The equivalence of GePUP-E with INSE
To derive GePUP-E from INSE, we split $\mathbf{u}$ into $\nabla\phi$ and $\mathbf{w}$ and suppress $\nabla\cdot\mathbf{w}$ via a heat equation.
Lemma 3.8
The GePUP-E formulation (32) is derived from the INSE in (1) by splitting the velocity $\mathbf{u}$ as $\mathbf{u}=\mathbf{w}+\nabla\phi$ and demanding that $\mathbf{w}$ satisfy (32b) and (37).
Proof
For an arbitrary vector field $\mathbf{u}$, (14) and (21) yield
$\Delta{\mathcal{P}}\mathbf{u}=\Delta(\mathbf{u}-\nabla\phi)=\Delta\mathbf{u}-\nabla\nabla\cdot\nabla\phi={\mathcal{B}}\mathbf{u}+\nabla\nabla\cdot{\mathcal{P}}\mathbf{u},$
and thus we have
$\Delta{\mathcal{P}}=\Delta-\nabla\nabla\cdot+\nabla\nabla\cdot{\mathcal{P}}.$ (40)
We deduce (32a) from
$\begin{array}[]{rl}\frac{\partial\mathbf{w}}{\partial t}=&\frac{\partial{\mathcal{P}}\mathbf{u}}{\partial t}={\mathcal{P}}\frac{\partial\mathbf{u}}{\partial t}={\mathcal{P}}\mathbf{a}-\nu(\Delta\mathbf{u}-\nabla\nabla\cdot\mathbf{u}+\nabla\nabla\cdot{\mathcal{P}}\mathbf{u}-\Delta{\mathcal{P}}\mathbf{u})\\ =&{\mathcal{P}}\mathbf{a}-\mathbf{a}^{*}+\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u}+\nu\nabla\nabla\cdot\mathbf{u}-\nu\nabla\nabla\cdot\mathbf{w}+\nu\Delta\mathbf{w}\\ =&\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u}-\nabla q+\nu\Delta\mathbf{w},\end{array}$ (41)
where the third step follows from (40), the fourth from (6), and the last from
$\nabla q:=\mathbf{a}^{*}-{\mathcal{P}}\mathbf{a}-\nu\nabla\nabla\cdot\mathbf{u}+\nu\nabla\nabla\cdot\mathbf{w};$ (42)
the above RHS is indeed a gradient because of (14) and (5).
(32e) follows from (37) and the divergence of (42), i.e.,
$\Delta q=\nabla\cdot\left(\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u}\right)-\left[\frac{\partial\nabla\cdot\mathbf{w}}{\partial t}-\nu(\Delta\nabla\cdot\mathbf{w})\right]\qquad\text{ in }\Omega.$
(32f) follows from Lemma 3.3 and the normal components of (32a) on $\partial\Omega$, i.e., $\mathbf{n}\cdot\nabla q=\mathbf{n}\cdot(\mathbf{g}-\mathbf{u}\cdot\nabla\mathbf{u}+\nu\Delta\mathbf{w})+\frac{\partial(\mathbf{n}\cdot\mathbf{w})}{\partial t}$. ∎
An alternative interpretation of (41) might be illuminating. A given scalar field $\phi$ furnishes a specific generic projection ${\mathcal{P}}_{\phi}\mathbf{u}=\mathbf{u}-\nabla\phi$ that perturbs $\mathbf{u}$ to be non-solenoidal. For any $\phi$, the composite operator ${\mathscr{P}}\circ{\mathcal{P}}_{\phi}$ is the identity on the divergence-free vector space, i.e., $\nabla\cdot\mathbf{u}=0$ implies ${\mathscr{P}}\circ{\mathcal{P}}_{\phi}\mathbf{u}={\mathscr{P}}\mathbf{w}=\mathbf{u}$ for any ${\mathcal{C}}^{1}$ scalar field $\phi$. If the evolution of $\mathbf{w}={\mathcal{P}}_{\phi}\mathbf{u}$ in (41) did not have the exponential decay of $\nabla\cdot\mathbf{w}$, we would have to apply $\mathscr{P}$} to $\frac{\partial\mathbf{w}}{\partial t}$ to recover the INSE. However, thanks to the exponential decay of $\nabla\cdot\mathbf{w}$ in Theorem 3.4, $\mathbf{w}$ converges to $\mathbf{u}$ in Lemma 3.7 and thus there is no need to recover $\mathbf{u}$ from $\mathbf{w}$.
Theorem 3.9
GePUP-E in Definition 3.1 is equivalent to the INSE in (1).
Proof
This follows directly from Lemma 3.5 and Lemma 3.8. ∎
3.4 The monotonic decrease of the total kinetic energy
Lemma 3.10
A vector field $\mathbf{v}\in\mathcal{C}^{1}(\Omega)$ with $\nabla\cdot\mathbf{v}=0$ in $\Omega$ and $\mathbf{v}\cdot\mathbf{n}=0$ on $\partial\Omega$ is orthogonal to the gradient field of any scalar function $\phi\in{\mathcal{C}}^{1}(\Omega)$, i.e.,
$\left\langle\mathbf{v},\nabla\phi\right\rangle=\int_{\Omega}\mathbf{v}\cdot\nabla\phi\,\mathrm{d}V=0.$ (43)
Proof
By the chain rule, we have
$\int_{\Omega}\mathbf{v}\cdot\nabla\phi\,\mathrm{d}V=\int_{\Omega}\nabla\cdot(\phi\mathbf{v})\mathrm{d}V-\int_{\Omega}\phi\nabla\cdot\mathbf{v}\,\mathrm{d}V=\int_{\partial\Omega}\phi\mathbf{v}\cdot\mathbf{n}\,\mathrm{d}A=0,$
where the second step follows from the divergence theorem and $\nabla\cdot\mathbf{v}=0$ in $\Omega$ and the last from $\mathbf{v}\cdot\mathbf{n}=0$ on $\partial\Omega$. ∎
As a benefit of Lemma 3.10, the accuracy of the computed velocity is largely decoupled from that of the pressure gradient; this orthogonality condition can be enforced to machine precision in the FV formulation.
Definition 3.11
The kinetic energy of a fluid with velocity $\mathbf{u}$ is
$E_{\text{kinetic}}:=\frac{1}{2}\|\mathbf{u}\|^{2}=\frac{1}{2}\int_{\Omega}\mathbf{u}\cdot\mathbf{u}\,\mathrm{d}V.$ (44)
Theorem 3.12
Suppose the body force $\mathbf{g}$ in (32a) is conservative, i.e., $\mathbf{g}=-\nabla\varphi$ for some scalar field $\varphi$ in $\Omega$. Then the evolution of the kinetic energy in the GePUP-E formulation in Definition 3.1 is governed by
$\frac{\mathrm{d}}{\mathrm{d}t}E_{\text{kinetic}}=-\nu\|\nabla\mathbf{u}\|^{2}:=-\nu\sum_{i=1}^{{\scriptsize\textup{D}}}\sum_{j=1}^{{\scriptsize\textup{D}}}\int_{\Omega}\left|\frac{\partial u_{i}}{\partial x_{j}}\right|^{2}\mathrm{d}V.$
Proof
Since $\varphi$ can be absorbed into $q$, the assumption $\mathbf{g}=\mathbf{0}$ has no loss of generality. The inner product of $\mathbf{u}$ and the momentum equation (32a) give
$(*):\quad\left\langle\frac{\partial\mathbf{w}}{\partial t},\mathbf{u}\right\rangle=-\left\langle\mathbf{u}\cdot\nabla\mathbf{u},\mathbf{u}\right\rangle-\left\langle\nabla q,\mathbf{u}\right\rangle+\left\langle\nu\Delta\mathbf{w},\mathbf{u}\right\rangle.$
The left-hand side (LHS) is computed as
$\begin{array}[]{rl}\left\langle\frac{\partial\mathbf{w}}{\partial t},\mathbf{u}\right\rangle&=\left\langle\frac{\partial(\mathbf{u}-\nabla\phi)}{\partial t},\mathbf{u}\right\rangle=\left\langle\frac{\partial\mathbf{u}}{\partial t},\mathbf{u}\right\rangle-\left\langle\nabla\frac{\partial\phi}{\partial t},\mathbf{u}\right\rangle=\left\langle\frac{\partial\mathbf{u}}{\partial t},\mathbf{u}\right\rangle\\ &=\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{1}{2}\int_{\Omega}\mathbf{u}\cdot\mathbf{u}\,\mathrm{d}V\right)=\frac{\mathrm{d}}{\mathrm{d}t}E_{\text{kinetic}},\end{array}$ (45)
where the second step follows from the commutativity of $\partial_{t}$ and $\nabla$, the third from (32c,d) and Lemma 3.10, and the last from (44).
The first RHS term in ()(*) vanishes because
$\begin{array}[]{rl}&\left\langle\mathbf{u}\cdot\nabla\mathbf{u},\mathbf{u}\right\rangle=\int_{\Omega}\mathbf{u}\cdot\left(\mathbf{u}\cdot\nabla\mathbf{u}\right)\mathrm{d}V=\int_{\Omega}u_{i}\left(u_{j}\frac{\partial u_{i}}{\partial x_{j}}\right)\mathrm{d}V=\frac{1}{2}\int_{\Omega}u_{j}\frac{\partial(u_{i}u_{i})}{\partial x_{j}}\mathrm{d}V\\ =&-\frac{1}{2}\int_{\Omega}u_{i}u_{i}\frac{\partial u_{j}}{\partial x_{j}}\mathrm{d}V+\frac{1}{2}\int_{\partial\Omega}u_{i}u_{i}u_{j}n_{j}\,\mathrm{d}A\\ =&-\frac{1}{2}\int_{\Omega}|\mathbf{u}|^{2}\left(\nabla\cdot\mathbf{u}\right)\mathrm{d}V+\frac{1}{2}\int_{\partial\Omega}|\mathbf{u}|^{2}\left(\mathbf{u}\cdot\mathbf{n}\right)\mathrm{d}A=0,\end{array}$
where the fourth step, in Einstein summation convention, follows from Lemma 2.6. The second term in ()(*) is also zero due to Lemma 3.10. The third term is
$\begin{array}[]{rl}&\left\langle\Delta\mathbf{w},\mathbf{u}\right\rangle=\left\langle\Delta(\mathbf{u}-\nabla\phi),\mathbf{u}\right\rangle=\left\langle\Delta\mathbf{u},\mathbf{u}\right\rangle-\left\langle\nabla\Delta\phi,\mathbf{u}\right\rangle=\left\langle\Delta\mathbf{u},\mathbf{u}\right\rangle\\ =&\int_{\Omega}u_{i}\Delta u_{i}\,\mathrm{d}V=-\int_{\Omega}\nabla u_{i}\cdot\nabla u_{i}\,\mathrm{d}V+\int_{\partial\Omega}u_{i}\frac{\partial u_{i}}{\partial\mathbf{n}}\mathrm{d}A=-\left\|\nabla\mathbf{u}\right\|^{2},\end{array}$ (46)
where the first step follows from the definition of $\mathbf{w}$, the second from the commutativity of $\Delta$ and $\nabla$, the third from (32c,d) and Lemma 3.10, the penultimate from Green’s formula (28), and the last from the no-slip condition of $\mathbf{u}$, which holds from Lemma 3.5 and (32b,d). ∎
3.5 Prominent features of the GePUP-E formulation
  1. (a)
    The sole evolutionary variable is the non-solenoidal velocity $\mathbf{w}$, with $\mathbf{u}$ determined from $\mathbf{w}$ via (32c,d) and $q$ from $\mathbf{u}$ and $\mathbf{w}$ via (32e,f). This chain of determination $\mathbf{w}\rightarrow\mathbf{u}\rightarrow q$ from Neumann BVPs is instantaneous and has nothing to do with time integration. Therefore, a time integrator in MOL can be employed in a black-box manner.
  2. (b)
    There is no ambiguity on which velocities should be projected and which should not in MOL; this resolves (TMR-1) discussed in Section 1.1.3.
  3. (c)
    Now that the main evolutionary variable $\mathbf{w}$ in (32) is formally non-solenoidal, the Leray-Helmholtz projection $\mathscr{P}$} only comes into the system (32) on the RHS. Although still contributing to the local truncation error, the approximation error of a discrete projection to $\mathscr{P}$} does not affect numerical stability of MOL; this resolves (TMR-2) in Section 1.1.3.
  4. (d)
    (32) comes with the built-in mechanisms of exponential decays of velocity divergence and total kinetic energy, which are conducive to the design of semi-discrete algorithms that ensure numerical stability and preserve physical structures of incompressible flows, c.f. Theorems 5.2 and 5.4.
4 The GePUP-ES formulation
The SAV approach, as introduced in shen18:SAV ; shen19:SAV_Review , has been proposed to develop time discretization schemes that are both efficient and stable for gradient flows. This approach was originally designed to create schemes that are linear, decoupled, unconditionally stable, and up to second-order accurate. It is also successfully extended to address the Navier-Stokes equations in li22:_new_sav_pressure_correction_inse ; lin19:_numer_appx_inses_auxiliary_energy_variable ; li20:_error_analysis_sav_mac .
More recently, based on the generalized SAV approach huang22:_new_class_imex_bdfk_sav , Huang et al. have devised high-order consistent splitting schemes for the Navier-Stokes equations, with periodic boundary conditions in huang21:_stability_error_analysis_higher_order_imex_ns_periodic_bc and no-slip boundary conditions in wu22:_new_class_higher_order_decoupled_inses_rotating .
In this section, we couple the GePUP-E formulation (32) to the SAV approach introduced in shen18:SAV ; shen19:SAV_Review to deal with the nonlinear convection term $\mathbf{u}\cdot\nabla\mathbf{u}$ so that unconditionally energy-stable numerical schemes can be constructed.
Definition 4.1
The GePUP-ES formulation of INSE on no-slip domains is
$\frac{\partial\mathbf{w}}{\partial t}$ $=\mathbf{g}-r(t)\mathbf{u}\cdot\nabla\mathbf{u}-\nabla q+\nu\Delta\mathbf{w}$ $\text{in }\Omega,$ (47a)
$\mathbf{w}\cdot\bm{\tau}$ $=0,\ \ \nabla\cdot\mathbf{w}=0\ \$$\text{on }\partial\Omega,$ (47b)
$\frac{\mathrm{d}r}{\mathrm{d}t}$ $=I_{cv}(\mathbf{u},\mathbf{u}),$ (47c)
$\mathbf{u}$ $=\mathscr{P}\mathbf{w}$ $\text{in }\Omega,$ (47d)
$\mathbf{u}\cdot\mathbf{n}$ $=0$ $\text{on }\partial\Omega,$ (47e)
$\Delta q$ $=\nabla\cdot\left(\mathbf{g}-r(t)\mathbf{u}\cdot\nabla\mathbf{u}\right)$ $\text{in }\Omega,$ (47f)
$\mathbf{n}\cdot\nabla q$ $=\mathbf{n}\cdot(\mathbf{g}-r(t)\mathbf{u}\cdot\nabla\mathbf{u}+\nu\Delta\mathbf{w})+\lambda\mathbf{n}\cdot\mathbf{w}$ $\text{on }\partial\Omega,$ (47g)
where $I_{cv}(\mathbf{u},\mathbf{v}):=\int_{\Omega}\left(\mathbf{u}\cdot\nabla\mathbf{u}\right)\cdot\mathbf{v}\,\mathrm{d}V$, the SAV $r(t)\equiv 1$, and $\lambda$ is a nonnegative penalty parameter.
The introduction of the SAV $r(t)\equiv 1$ immediately implies $\frac{\mathrm{d}r}{\mathrm{d}t}=0$. We can define the evolution of $r(t)$ as the ODE in (47c) because, for no-slip conditions, we always have $\left\langle\mathbf{u}\cdot\nabla\mathbf{u},\mathbf{u}\right\rangle=0$. As a newly added evolutionary variable, the SAV $r(t)$ is a double-edged sword. On the one hand, it leads to a tighter coupling between $\mathbf{u}$ and $\mathbf{w}$, which makes it difficult to orchestrate an implicit or semi-implicit RK method as solving a sequence of linear systems; see the discussions in Section 5.2. On the other hand, it preserves the monotonic decrease of the kinetic energy; see Theorems 4.2 and 5.4.
Theorem 4.2
Suppose that the body force $\mathbf{g}$ is conservative. Then the energy dissipation of the GePUP-ES formulation (47) is governed by
$\frac{\mathrm{d}}{\mathrm{d}t}\left(E_{\mathrm{kinetic}}+\frac{r^{2}}{2}\right)=-\nu\|\nabla\mathbf{u}\|^{2}.$
Proof
Take inner product with $\mathbf{u}$ in (47a), multiply (47c) by $r(t)$, add up the resulting two equations, and we cancel the integral of the convection term to obtain $\left\langle\frac{\partial\mathbf{w}}{\partial t},\mathbf{u}\right\rangle+r(t)r^{\prime}(t)=\left\langle\mathbf{g}-\nabla q+\nu\Delta\mathbf{w},\mathbf{u}\right\rangle$. The rest of the proof follows from (45), (46), and Lemma 3.10. ∎
Theorem 3.4 also holds for GePUP-ES with exactly the same proof. Similarly, the GePUP-ES formulation (47) retains the advantage of GePUP-E that the temporal integration and spatial discretization are completely decoupled. Hence the fourth-order finite-volume discrete operators in zhang16:_GePUP can be reused and we will focus on temporal integration hereafter.
5 Algorithms
Based on the GePUP-ES formulation, we construct numerical algorithms to preserve the monotonic decrease of the modified kinetic energy, the divergence residue, and the magnitude of the normal velocity on the domain boundary.
5.1 Semi-discrete GePUP-ES-RK schemes
These schemes follow directly from discretizing the GePUP-ES formulation (47) in time by the RK method (29).
Definition 5.1
A GePUP-ES-RK scheme for solving INSE with no-slip conditions is a semidiscrete algorithm of the form
$\left\{\begin{array}[]{rlcl}\mathbf{w}^{n+1}&=\mathbf{u}^{n}+k\sum_{i=1}^{s}b_{i}\bm{\rho}^{(i)}&&\text{in }\Omega,\\ r^{n+1}&=r^{n}+k\sum_{i=1}^{s}b_{i}I_{cv}(\widetilde{\mathbf{u}}^{(i)},\mathbf{u}^{(i)}),&&\\ \mathbf{u}^{n+1}&=\mathscr{P}\mathbf{w}^{n+1}&&\text{in }\Omega,\\ \mathbf{n}\cdot\mathbf{u}^{n+1}&=0&&\text{on }\partial\Omega,\end{array}\right$. (48)
where $s$ is the number of stages of the employed RK method $(A,\mathbf{b},\mathbf{c})$, the integral $I_{cv}(\widetilde{\mathbf{u}}^{(i)},\mathbf{u}^{(i)})$ is the same as that in Definition 4.1, the auxiliary velocity $\widetilde{\mathbf{u}}^{(i)}$ is a suitable explicit approximation to $\mathbf{u}(t^{n}+c_{i}k)$ and
$\Delta q^{(i)}$ $=\nabla\cdot\left(\mathbf{g}^{(i)}-r^{(i)}\widetilde{\mathbf{u}}^{(i)}\cdot\nabla\widetilde{\mathbf{u}}^{(i)}\right)$ $\text{in }\Omega,$ (49a)
$\mathbf{n}\cdot\nabla q^{(i)}$ $=\mathbf{n}\cdot\left(\mathbf{g}^{(i)}-r^{(i)}\widetilde{\mathbf{u}}^{(i)}\cdot\nabla\widetilde{\mathbf{u}}^{(i)}+\nu\Delta\mathbf{w}^{(i)}\right)+\lambda\mathbf{n}\cdot\mathbf{w}^{(i)}$ $\text{on }\partial\Omega,$ (49b)
$\bm{\rho}^{(i)}$ $:=\mathbf{g}^{(i)}-r^{(i)}\widetilde{\mathbf{u}}^{(i)}\cdot\nabla\widetilde{\mathbf{u}}^{(i)}-\nabla q^{(i)}+\nu\Delta\mathbf{w}^{(i)},$ (49c)
$\mathbf{w}^{(i)}$ $=\mathbf{u}^{n}+k\sum_{j=1}^{s}a_{i,j}\bm{\rho}^{(j)}$ $\text{in }\Omega,$ (49d)
$\mathbf{w}^{(i)}\cdot\bm{\tau}$ $=0,~\nabla\cdot\mathbf{w}^{(i)}=0$ $\text{on }\partial\Omega,$ (49e)
$r^{(i)}$ $=r^{n}+k\sum_{j=1}^{s}a_{i,j}I_{cv}(\widetilde{\mathbf{u}}^{(j)},\mathbf{u}^{(j)}),$ (49f)
$\mathbf{u}^{(i)}$ $=\mathscr{P}\mathbf{w}^{(i)}$ $\text{in }\Omega,$ (49g)
$\mathbf{n}\cdot\mathbf{u}^{(i)}$ $=0$ $\text{on }\partial\Omega.$ (49h)
As suggested by (49d) and the first equation of (48), $\mathbf{u}^{n}$ is used as the initial condition of $\mathbf{w}$ for time integration within the interval $t^{n},t^{n+1}]$. However, it is emphasized that we do not write $\mathbf{w}^{n}=\mathbf{u}^{n}$ because this would cause a notation clash to $\mathbf{w}^{n+1}$, which, by (49), is the evolutionary velocity before the final projection and thus needs not be divergence free. In Definition 5.1, the approximations $\widetilde{\mathbf{u}}^{(i)}$ in $\Omega$ may be obtained by polynomial interpolation based on stage values of recent time steps. In this work, we fit a cubic polynomial $\mathbf{p}$ from the known velocity $\hat{\mathbf{u}}_{n}^{(j)}$ at time instances $t^{n}+\hat{c}_{j}k$ with $j=0,1,2,3$ and then approximate $\mathbf{u}(t^{n}+c_{i}k)$ with
$\forall i=1,\ldots,s,\qquad\widetilde{\mathbf{u}}_{n}^{(i)}=\mathbf{p}\left(c_{i}k\right):=\sum_{j=0}^{3}\hat{\mathbf{u}}_{n}^{(j)}\prod\limits_{\ell\neq j}\frac{c_{i}-\hat{c}_{\ell}}{\hat{c}_{j}-\hat{c}_{\ell}}$; (50)
  1. (EAV-1)
    for the first two time steps $n=0,1$, we calculate $\hat{\mathbf{u}}^{(j)}$ by GePUP-ERK zhang16:_GePUP , the explicit RK method for solving the GePUP formulation, with time step size $\frac{1}{3}k$ and then fit $\mathbf{p}$ in (50) with $\left[\hat{c}_{0},\hat{c}_{1},\hat{c}_{2},\hat{c}_{3}\right]=\left[0,\frac{1}{3},\frac{2}{3},1\right]$;
  2. (EAV-2)
    for $n\geq 2$, we first calculate $\hat{\mathbf{u}}^{(3)}=\mathbf{u}^{n+1}$ by GePUP-ERK with time step size $k$ and then fit $\mathbf{p}$ in (50) with $\left[\hat{c}_{0},\hat{c}_{1},\hat{c}_{2},\hat{c}_{3}\right]=\left[-2,-1,0,1\right]$.
We emphasize that $\hat{\mathbf{u}}_{n}^{(j)}$, $\widetilde{\mathbf{u}}_{n}^{(i)}$, and $\mathbf{u}(t^{n}+c_{i}k)$ are all at the same location so that the interpolation is only in time; see Figure 1 for an illustration.
$t^{n}$$t^{n+1}$$\widetilde{\mathbf{u}}_{n}^{\left(1\right)}$$\widetilde{\mathbf{u}}_{n}^{\left(2\right)}$$\widetilde{\mathbf{u}}_{n}^{\left(3\right)}$$t^{n}+c_{1}k$$t^{n}+c_{2}k$$t^{n}+c_{3}k$$t^{n}$$t^{n}+\frac{1}{3}k$$t^{n}+\frac{3}{3}k=t^{n+1}$$\mathbf{u}^{n}$$\mathbf{u}^{n+1}$$\hat{\mathbf{u}}_{n}^{\left(0\right)}=\mathbf{u}^{n}$$\hat{\mathbf{u}}_{n}^{\left(1\right)}$$\hat{\mathbf{u}}_{n}^{\left(3\right)}$$t^{n}+\frac{2}{3}k$$\hat{\mathbf{u}}_{n}^{\left(2\right)}$
(a) (EAV-1) for $n=0,1$
$t^{n}$$t^{n+1}$$t^{n-2}$$\widetilde{\mathbf{u}}_{n}^{\left(1\right)}$$\widetilde{\mathbf{u}}_{n}^{\left(2\right)}$$\widetilde{\mathbf{u}}_{n}^{\left(3\right)}$$t^{n}+c_{1}k$$t^{n}+c_{2}k$$t^{n}+c_{3}k$$t^{n}$$t^{n-1}$$t^{n+1}=t^{n}+k$$\mathbf{u}^{n}$$\mathbf{u}^{n+1}$$\hat{\mathbf{u}}_{n}^{\left(2\right)}=\mathbf{u}^{n}$$\hat{\mathbf{u}}_{n}^{\left(1\right)}=\mathbf{u}^{n-1}$$\hat{\mathbf{u}}_{n}^{\left(3\right)}$$\hat{\mathbf{u}}_{n}^{\left(0\right)}=\mathbf{u}^{n-2}$
(b) (EAV-2) for $n\geq 2$
Figure 1: Estimating the auxiliary velocity $\widetilde{\mathbf{u}}^{(i)}$ (EAV) in GePUP-ES-RK by (50).
Theorem 5.2
Suppose the RK method employed in the GePUP-ES-RK scheme (48) is algebraically stable in the sense of Definition 2.9. Then (34) implies
$\forall n\in\mathbb{N},\quad\left\|\nabla\cdot\mathbf{w}^{n+1}\right\|^{2}-\left\|\nabla\cdot\mathbf{w}^{n}\right\|^{2}\leq-2k\nu\sum_{i=1}^{s}b_{i}\left\|\nabla\left(\nabla\cdot\mathbf{w}^{(i)}\right)\right\|^{2},$ (51)
$\forall n\!\in\!\mathbb{N},\ \forall\mathbf{x}\!\in\!\partial\Omega,~\left|\mathbf{n}\!\cdot\!\mathbf{w}^{n+1}(\mathbf{x})\right|^{2}\!-\!\left|\mathbf{n}\!\cdot\!\mathbf{w}^{n}(\mathbf{x})\right|^{2}\!\leq\!-2k\lambda\sum_{i=1}^{s}b_{i}\left|\mathbf{n}\!\cdot\!\mathbf{w}^{(i)}(\mathbf{x})\right|^{2}\!.$ (52)
Proof
Take divergence of (49c), apply (49a) and the commutativity of $\Delta$ and $\nabla\cdot$, and we have $\nabla\cdot\bm{\rho}^{(i)}=\nu\Delta\left(\nabla\cdot\mathbf{w}^{(i)}\right)$. Then
$\begin{array}[]{rl}&\left\langle\nabla\cdot\mathbf{w}^{n+1},\nabla\cdot\mathbf{w}^{n+1}\right\rangle\\ =&\left\langle\nabla\cdot\mathbf{u}^{n}+k\sum_{i=1}^{s}b_{i}\nabla\cdot\bm{\rho}^{(i)},\nabla\cdot\mathbf{u}^{n}+k\sum_{j=1}^{s}b_{j}\nabla\cdot\bm{\rho}^{(j)}\right\rangle\\ =&\left\langle\nabla\cdot\mathbf{u}^{n},\nabla\cdot\mathbf{u}^{n}\right\rangle+k\sum_{i=1}^{s}b_{i}\left\langle\nabla\cdot\mathbf{u}^{n},\nabla\cdot\bm{\rho}^{(i)}\right\rangle\\ &+k\sum_{j=1}^{s}b_{j}\left\langle\nabla\cdot\mathbf{u}^{n},\nabla\cdot\bm{\rho}^{(j)}\right\rangle+k^{2}\sum_{i=1}^{s}\sum_{j=1}^{s}b_{i}b_{j}\left\langle\nabla\cdot\bm{\rho}^{(i)},\nabla\cdot\bm{\rho}^{(j)}\right\rangle\\ =&\left\langle\nabla\cdot\mathbf{u}^{n},\nabla\cdot\mathbf{u}^{n}\right\rangle+k^{2}\sum_{i=1}^{s}\sum_{j=1}^{s}b_{i}b_{j}\left\langle\nabla\cdot\bm{\rho}^{(i)},\nabla\cdot\bm{\rho}^{(j)}\right\rangle\\ &+k\sum_{i=1}^{s}b_{i}\left\langle\nabla\cdot\mathbf{w}^{(i)}-k\sum_{j=1}^{s}a_{i,j}\nabla\cdot\bm{\rho}^{(j)},\nabla\cdot\bm{\rho}^{(i)}\right\rangle\\ &+k\sum_{j=1}^{s}b_{j}\left\langle\nabla\cdot\mathbf{w}^{(j)}-k\sum_{i=1}^{s}a_{j,i}\nabla\cdot\bm{\rho}^{(i)},\nabla\cdot\bm{\rho}^{(j)}\right\rangle\\ =&\left\langle\nabla\cdot\mathbf{u}^{n},\nabla\cdot\mathbf{u}^{n}\right\rangle-2\nu k\sum_{i=1}^{s}b_{i}\left\|\nabla\left(\nabla\cdot\mathbf{w}^{(i)}\right)\right\|^{2}\\ &-k^{2}\sum_{i=1}^{s}\sum_{j=1}^{s}m_{i,j}\left\langle\nabla\cdot\bm{\rho}^{(i)},\nabla\cdot\bm{\rho}^{(j)}\right\rangle,\end{array}$
where the last step follows from (31), (28), and (49e). By Definition 2.9, the algebraic stability implies the symmetric positive semi-definiteness of $M$ and thus we have $M=O\Lambda O^{T}$, where $O$ is an orthogonal matrix and $\Lambda$ is a diagonal matrix with $\lambda_{\ell}=\Lambda_{\ell,\ell}\geq 0$ for each $\ell=1,\ldots,s$. Hence (51) follows from
$\begin{array}[]{rl}&\sum_{i=1}^{s}\sum_{j=1}^{s}m_{i,j}\left\langle\nabla\cdot\bm{\rho}^{(i)},\nabla\cdot\bm{\rho}^{(j)}\right\rangle\\ =&\sum_{i=1}^{s}\sum_{j=1}^{s}\left(\sum_{\ell=1}^{s}o_{i,\ell}\lambda_{\ell}o_{j,\ell}\right)\left\langle\nabla\cdot\bm{\rho}^{(i)},\nabla\cdot\bm{\rho}^{(j)}\right\rangle\\ =&\sum_{\ell=1}^{s}\lambda_{\ell}\left\langle\sum_{i=1}^{s}o_{i,\ell}\nabla\cdot\bm{\rho}^{(i)},\sum_{j=1}^{s}o_{j,\ell}\nabla\cdot\bm{\rho}^{(j)}\right\rangle\\ =&\sum_{\ell=1}^{s}\lambda_{\ell}\left\|\sum_{i=1}^{s}o_{i,\ell}\nabla\cdot\bm{\rho}^{(i)}\right\|^{2}\geq 0\end{array}$
and the fact that the Leray-Helmholtz projection $\mathscr{P}$} do not increase the divergence of velocity, i.e., $\|\nabla\cdot\mathbf{u}^{n}\|^{2}=\|\nabla\cdot{\mathscr{P}}\mathbf{w}^{n}\|^{2}\leq\|\nabla\cdot\mathbf{w}^{n}\|^{2}$.
(49b) and the normal component of (49c) imply $\mathbf{n}\cdot\bm{\rho}^{(i)}=-\lambda\mathbf{n}\cdot\mathbf{w}^{(i)}$. Then (52) follows from arguments similar to those for (51) and
$\forall n\in\mathbb{N},\ \forall\mathbf{x}\in\partial\Omega,\ \left|\mathbf{n}\cdot\mathbf{u}^{n}(\mathbf{x})\right|^{2}-\left|\mathbf{n}\cdot\mathbf{w}^{n}(\mathbf{x})\right|^{2}=-\left|\mathbf{n}\cdot\mathbf{w}^{n}(\mathbf{x})\right|^{2}\leq 0.\hskip 22.76219pt\qed$
In the fully discrete sense, we discretize each continuous operator in (51) add the corresponding discretization error, and obtain a discrete version of (51). As discussed in Section 6, this discrete version of (51) is helpful in understanding the evolution of the discrete velocity divergence.
Interestingly, (51) is useless for first-order finite difference/volume methods. Suppose $\mathbf{w}^{n}=\mathbf{w}(t^{n})+O(h^{2})$. Then, due to the factor $k$ and the second derivative on the RHS of (51), the discrete version of (51) has an $O(1)$ error. In contrast, a fourth-order method has $\mathbf{w}^{n}=\mathbf{w}(t^{n})+O(h^{4})$. With the added discretization error approaching zero as $h,k\rightarrow 0$, the discrete version of (51) indeed has control over the evolution of the discrete velocity divergence. The above discussion suggests a crucial advantage of fourth-order methods over first-order methods.
Corollary 5.3
Suppose the RK method employed in the GePUP-ES-RK scheme (48) is algebraically stable in the sense of Definition 2.9. Then the initial condition (33) implies
$\forall n\in\mathbb{N}^{+},$ $\quad\left\{\begin{array}[]{l}\left\|\nabla\cdot\mathbf{w}^{n}\right\|=0;\\ \forall i=1,\ldots,s,\quad\nabla\cdot\mathbf{w}^{(i)}=0,\end{array}\right$. (55)
$\forall n\in\mathbb{N}^{+},$ $\quad\left\{\begin{array}[]{l}\mathbf{n}\cdot\mathbf{w}^{n}|_{\partial\Omega}=0;\\ \forall i=1,\ldots,s,\quad\mathbf{n}\cdot\mathbf{w}^{(i)}|_{\partial\Omega}=0.\end{array}\right$. (58)
Proof
(33) gives $\nabla\cdot\mathbf{w}(t_{0})=\nabla\cdot\mathbf{u}(t_{0})=0$ and $\mathbf{n}\cdot\mathbf{w}(t_{0})|_{\partial\Omega}=\mathbf{n}\cdot\mathbf{u}(t_{0})|_{\partial\Omega}=0$. Then the first clauses of (55) and (58) follow from Theorem 5.2 and the first condition in Definition 2.9. (51) dictates $\|\nabla(\nabla\cdot\mathbf{w}^{(i)})\|=0$ for each stage. Then the second clause of (55) follows from the boundary condition $\nabla\cdot\mathbf{w}|_{\partial\Omega}=0$ in (32b). Similarly, (52) dictates the second clause of (58). ∎
Of course Corollary 5.3 holds only in the semi-discrete sense. Since Corollary 5.3 is used in proving Theorem 5.4, the decay of the modified energy in the fully discrete sense depends on that of the discrete velocity divergence.
Theorem 5.4
Suppose that the body force $\mathbf{g}$ in the GePUP-E formulation is conservative, that the initial condition of $\mathbf{w}$ is (33), and that the employed RK method in (48) is algebraically stable in the sense of Definition 2.9. Then the energy dissipation of the GePUP-ES-RK scheme (48) is governed by
$\mathcal{E}\left(t^{n+1}\right)-\mathcal{E}(t^{n})\leq-k\nu\sum_{i=1}^{s}b_{i}\left\|\nabla\mathbf{u}^{(i)}\right\|^{2},$ (59)
where the modified energy is defined as $\mathcal{E}(t^{n}):=\frac{1}{2}\left(\left\|\mathbf{u}^{n}\right\|^{2}+\left|r^{n}\right|^{2}\right)$.
Proof
Denote $\bm{\sigma}^{(i)}:=\mathscr{P}\bm{\rho}^{(i)}$ and we have from (49c,d) and (48)
$\begin{array}[]{c}\forall i=1,\ldots,s,\ \ \mathbf{u}^{(i)}=\mathscr{P}\mathbf{w}^{(i)}=\mathscr{P}\left(\mathbf{u}^{n}+k\sum_{j=1}^{s}a_{i,j}\bm{\rho}^{(j)}\right)=\mathbf{u}^{n}+k\sum_{j=1}^{s}a_{i,j}\bm{\sigma}^{(j)};\\ \mathbf{u}^{n+1}=\mathscr{P}\mathbf{w}^{n+1}=\mathscr{P}\left(\mathbf{u}^{n}+k\sum_{i=1}^{s}b_{i}\bm{\rho}^{(i)}\right)=\mathbf{u}^{n}+k\sum_{i=1}^{s}b_{i}\bm{\sigma}^{(i)}.\end{array}$
It follows that
$\begin{array}[]{rl}&\left\langle\mathbf{u}^{n+1},\mathbf{u}^{n+1}\right\rangle=\left\langle\mathbf{u}^{n}+k\sum_{i=1}^{s}b_{i}\bm{\sigma}^{(i)},\mathbf{u}^{n}+k\sum_{j=1}^{s}b_{j}\bm{\sigma}^{(j)}\right\rangle\\ =&\left\langle\mathbf{u}^{n},\mathbf{u}^{n}\right\rangle+k\sum_{i=1}^{s}b_{i}\left\langle\mathbf{u}^{n},\bm{\sigma}^{(i)}\right\rangle+k\sum_{j=1}^{s}b_{j}\left\langle\mathbf{u}^{n},\bm{\sigma}^{(j)}\right\rangle\\ &+k^{2}\sum_{i=1}^{s}\sum_{j=1}^{s}b_{i}b_{j}\left\langle\bm{\sigma}^{(i)},\bm{\sigma}^{(j)}\right\rangle\\ =&\left\langle\mathbf{u}^{n},\mathbf{u}^{n}\right\rangle+k\sum_{i=1}^{s}b_{i}\left\langle\mathbf{u}^{(i)}-k\sum_{j=1}^{s}a_{i,j}\bm{\sigma}^{(j)},\bm{\sigma}^{(i)}\right\rangle\\ &+k\sum_{j=1}^{s}b_{j}\left\langle\mathbf{u}^{(j)}-k\sum_{i=1}^{s}a_{j,i}\bm{\sigma}^{(i)},\bm{\sigma}^{(j)}\right\rangle+k^{2}\sum_{i=1}^{s}\sum_{j=1}^{s}b_{i}b_{j}\left\langle\bm{\sigma}^{(i)},\bm{\sigma}^{(j)}\right\rangle\\ =&\left\langle\mathbf{u}^{n},\mathbf{u}^{n}\right\rangle+2k\sum_{i=1}^{s}b_{i}\left\langle\mathbf{u}^{(i)},\bm{\sigma}^{(i)}\right\rangle-k^{2}\sum_{i=1}^{s}\sum_{j=1}^{s}m_{i,j}\left\langle\bm{\sigma}^{(i)},\bm{\sigma}^{(j)}\right\rangle\end{array}$
where the last step is due to (31). The second condition in Definition 2.9 gives
$\begin{array}[]{l}(*):\quad\left\langle\mathbf{u}^{n+1},\mathbf{u}^{n+1}\right\rangle\leq\left\langle\mathbf{u}^{n},\mathbf{u}^{n}\right\rangle+2k\sum_{i=1}^{s}b_{i}\left\langle\mathbf{u}^{(i)},\bm{\sigma}^{(i)}\right\rangle.\end{array}$
Write $\alpha_{i}:=I_{cv}(\widetilde{\mathbf{u}}^{(i)},\mathbf{u}^{(i)})=\left\langle\widetilde{\mathbf{u}}^{(i)}\cdot\nabla\widetilde{\mathbf{u}}^{(i)},\mathbf{u}^{(i)}\right\rangle$ for each $i=1,\ldots,s$. Then
$\begin{array}[]{rl}&\left\langle\bm{\sigma}^{(i)},\mathbf{u}^{(i)}\right\rangle=\left\langle\mathscr{P}\bm{\rho}^{(i)},\mathbf{u}^{(i)}\right\rangle=\left\langle\bm{\rho}^{(i)},\mathbf{u}^{(i)}\right\rangle\\ =&\left\langle\mathbf{g}^{(i)}-\nabla q^{(i)},\mathbf{u}^{(i)}\right\rangle-r^{(i)}\alpha_{i}+\nu\left\langle\Delta\mathbf{w}^{(i)},\mathbf{u}^{(i)}\right\rangle=-r^{(i)}\alpha_{i}-\nu\left\|\nabla\mathbf{u}^{(i)}\right\|^{2},\end{array}$
where the first equality follows from $\bm{\sigma}^{(i)}=\mathscr{P}\bm{\rho}^{(i)}$, the second from (19), (49g,h), and Lemma 3.10, the third from (49c) and the definition of $\alpha_{i}$, and the last from $\mathbf{g}$ being conservative, Lemma 3.10, (46), and Corollary 5.3.
Since the Leray-Helmholtz projection $\mathscr{P}$ has no control over the tangential velocity, in (49g) we would have $\bm{\tau}\cdot\mathbf{u}^{(i)}\neq 0$ if $\nabla\cdot\mathbf{w}^{(i)}\neq 0$. Fortunately Corollary 5.3 dictates $\nabla\cdot\mathbf{w}^{(i)}=0$ and $\mathbf{n}\cdot\mathbf{w}^{(i)}|_{\partial\Omega}=0$, then in (49g) $\mathscr{P}$ reduces to the identity. Thus the second integral in the last line of (46) vanishes in GePUP-ES-RK. A related observation is that, although $\nabla q$ is orthogonal to $\mathbf{w}$, it cannot be arbitrary as it must make $\bm{\rho}^{(j)}$ in (49d) divergence-free; otherwise in (49g) $\mathscr{P}$ would not reduce to the identity.
Substitute $\left\langle\mathbf{u}^{(i)},\bm{\sigma}^{(i)}\right\rangle=-r^{(i)}\alpha_{i}-\nu\left\|\nabla\mathbf{u}^{(i)}\right\|^{2}$ into ()(*) and we have
$\begin{array}[]{l}\left\langle\mathbf{u}^{n+1},\mathbf{u}^{n+1}\right\rangle\leq\left\langle\mathbf{u}^{n},\mathbf{u}^{n}\right\rangle-2k\sum_{i=1}^{s}b_{i}\alpha_{i}r^{(i)}-2\nu k\sum_{i=1}^{s}b_{i}\left\|\nabla\mathbf{u}^{(i)}\right\|^{2}.\end{array}$
Similarly, the positive semi-definiteness of $M$ in Definition 2.9 yields
$\begin{array}[]{rl}&\left|r^{n+1}\right|^{2}=\left(r^{n}+k\sum_{i=1}^{s}b_{i}\alpha_{i}\right)^{2}\\ =&\left|r^{n}\right|^{2}+k\sum_{i=1}^{s}b_{i}\alpha_{i}r^{n}+k\sum_{j=1}^{s}b_{j}\alpha_{j}r^{n}+k^{2}\sum_{i=1}^{s}\sum_{j=1}^{s}b_{i}b_{j}\alpha_{i}\alpha_{j}\\ =&\left|r^{n}\right|^{2}+k\sum_{i=1}^{s}b_{i}\alpha_{i}\left(r^{(i)}-k\sum_{j=1}^{s}a_{i,j}\alpha_{j}\right)\\ &+k\sum_{j=1}^{s}b_{j}\alpha_{j}\left(r^{(j)}-k\sum_{i=1}^{s}a_{j,i}\alpha_{i}\right)+k^{2}\sum_{i=1}^{s}\sum_{j=1}^{s}b_{i}b_{j}\alpha_{i}\alpha_{j}\\ =&\left|r^{n}\right|^{2}+2k\sum_{i=1}^{s}b_{i}\alpha_{i}r^{(i)}-k^{2}\sum_{i=1}^{s}\sum_{j=1}^{s}m_{i,j}\alpha_{i}\alpha_{j}\\ \leq&\left|r^{n}\right|^{2}+2k\sum_{i=1}^{s}b_{i}\alpha_{i}r^{(i)}.\end{array}$
The proof is completed by summing up the above two inequalities. ∎
In the last step of the above proof, it is the auxiliary variable $r$ in the GePUP-ES formulation that leads to the cancellation of $2k\sum_{i=1}^{s}b_{i}\alpha_{i}r^{(i)}$.
Except on staggered grids, the initial condition of $\mathbf{w}$ being incompressible in (33) cannot be exactly fulfilled in practical computations. However, the mechanism of divergence decay ensures that $\mathbf{w}$ converge to $\mathbf{u}$ sufficiently fast; see Theorem 5.3. This is also related to the millennium problem on the well-posedness of the INSE. If the solution of the INSE blows up, then its divergence must blow up first. Therefore, our computation only works when the INSE admits a bounded solution. After all, one cannot expect to solve the millennium problem by reformulating INSE.
5.2 Semi-implicit GePUP-ES-SDIRK schemes
In light of Theorems 5.2 and 5.4, one way to preserve the monotonic decrease of the kinetic energy and the exponential decay of the divergence residue is to employ an algebraically stable RK method in GePUP-ES-RK. Gauss-Legendre RK methods are algebraically stable and have a minimal number of stages for a given temporal order of accuracy. However, their employment in GePUP-ES-RK necessitates either the coupling of all intermediate stage values of $\mathbf{w}$ or the use of complex arithmetic. Thus we turn to singly diagonal implicit RK (SDIRK) methods that satisfy
$a_{i,j}=\begin{cases}0&\text{if }i<j;\\ \gamma\neq 0&\text{if }i=j,\end{cases}$ (60)
aiming to design a family of GePUP-ES-SDIRK schemes that only consist of solving a sequence of linear BVPs with real arithmetic, one intermediate stage at a time. The core difficulty for this, as mentioned in Section 4, is the nonlinear tight coupling of $\mathbf{u}$, $\mathbf{w}$, $q$ and $r$. Our solution is
Definition 5.5
A GePUP-ES-SDIRK (GES) scheme is a GePUP-ES-RK scheme where an algebraically stable SDIRK method is employed as the RK method and stage values for each intermediate stage $i=1,\ldots,s$ are decomposed as
$\left\{\begin{array}[]{rl}\mathbf{w}^{\left(i\right)}&=\mathbf{w}_{0}^{(i)}+\sum_{j=1}^{i}r^{(j)}\mathbf{w}_{j}^{(i)},\\ \mathbf{u}^{(i)}&=\mathbf{u}_{0}^{(i)}+\sum_{j=1}^{i}r^{(j)}\mathbf{u}_{j}^{(i)},\\ q^{(i)}&=q_{1}^{(i)}+r^{(i)}q_{2}^{(i)},\end{array}\right$. (61)
where $\mathbf{w}_{0}^{(i)},\mathbf{w}_{1}^{(i)},\ldots,\mathbf{w}_{i}^{(i)}$, $\mathbf{u}_{0}^{(i)},\mathbf{u}_{1}^{(i)},\ldots,\mathbf{u}_{i}^{(i)}$, and $q_{1}^{(i)},q_{2}^{(i)}$ are auxiliary variables for the $i$th stage, which consists of steps as follows,
  1. (GES.1)
    solve for $q_{1}^{(i)}$ from the Neumann BVP
    $\left\{\begin{array}[]{rll}\Delta q_{1}^{(i)}&=\nabla\cdot\mathbf{g}^{(i)}&\text{in }\Omega,\\ \mathbf{n}\cdot\nabla q_{1}^{(i)}&=\mathbf{n}\cdot\left(\mathbf{g}^{(i)}+\nu\Delta\widetilde{\mathbf{w}}^{(i)}\right)+\lambda\mathbf{n}\cdot\widetilde{\mathbf{w}}^{(i)}&\text{on }\partial\Omega,\end{array}\right$.
    where $\widetilde{\mathbf{w}}^{(i)}$ is an approximation of $\mathbf{w}^{(i)}$ obtained by (50) and (EAV-1,2).
  2. (GES.2)
    solve for $q_{2}^{(i)}$ from the Neumann BVP
    $\left\{\begin{array}[]{rll}\Delta q_{2}^{(i)}&=-\nabla\cdot\left(\widetilde{\mathbf{u}}^{(i)}\cdot\nabla\widetilde{\mathbf{u}}^{(i)}\right)&\text{in }\Omega,\\ \mathbf{n}\cdot\nabla q_{2}^{(i)}&=-\mathbf{n}\cdot\left(\widetilde{\mathbf{u}}^{(i)}\cdot\nabla\widetilde{\mathbf{u}}^{(i)}\right)&\text{on }\partial\Omega.\end{array}\right$.
  3. (GES.3)
    solve for $\mathbf{w}_{0}^{(i)}$ and $\mathbf{w}_{\ell}^{(i)}$ where $\ell=1,\ldots,i$ from BVPs with their boundary conditions in (49e).
    $\left\{\begin{array}[]{rl}(1-\nu\gamma k\Delta)\mathbf{w}_{0}^{(i)}\par&=\mathbf{u}^{n}+k\sum\limits_{j=1}^{i}a_{i,j}\left(\mathbf{g}^{(j)}-\nabla q_{1}^{(j)}\right)+\nu k\sum\limits_{j=1}^{i-1}a_{i,j}\Delta\mathbf{w}_{0}^{(j)};\\ (1-\nu\gamma k\Delta)\mathbf{w}_{\ell}^{(i)}&=-ka_{i,\ell}\left(\widetilde{\mathbf{u}}^{(\ell)}\cdot\nabla\widetilde{\mathbf{u}}^{(\ell)}+\nabla q_{2}^{(\ell)}\right)+\nu k\sum\limits_{j=\ell}^{i-1}a_{i,j}\Delta\mathbf{w}_{\ell}^{(j)}.\end{array}\right$.
  4. (GES.4)
    compute $\mathbf{u}_{\ell}^{(i)}={\mathscr{P}}\mathbf{w}_{\ell}^{(i)}$ for $\ell=0,1,\ldots,i$.
  5. (GES.5)
    define $\overline{\mathbf{u}}^{(i)}:=\mathbf{u}_{0}^{(i)}+\sum_{\ell=1}^{i-1}r^{(\ell)}\mathbf{u}_{\ell}^{(i)}$ and calculate $r^{(i)}$ by
    $\left(1-\gamma kI_{cv}\left(\widetilde{\mathbf{u}}^{(i)},\mathbf{u}^{(i)}_{i}\right)\right)r^{(i)}\!=\!r^{n}+\gamma kI_{cv}\left(\widetilde{\mathbf{u}}^{(i)}\!,\!\overline{\mathbf{u}}^{(i)}\right)+k\sum_{j=1}^{i-1}a_{i,j}I_{cv}\left(\widetilde{\mathbf{u}}^{(j)}\!,\!{\mathbf{u}}^{(j)}\right).$
  6. (GES.6)
    calculate $\mathbf{w}^{(i)}$ and $\mathbf{u}^{(i)}$ by (61).
The above steps (GES.1–6) are direct consequences of the GePUP-ES-RK scheme (48), the property (60) of SDIRK, and the decomposition (61).
Substitute the third decomposition in (61) into (49a,b), separate the terms with and without $r^{(i)}$, notice that $r^{(i)}$ is a scalar, and we have (GES.1–2). In step (GES.1), we decouple $q^{(i)}$ from $\mathbf{w}^{(i)}$ by approximating $\mathbf{w}^{(i)}$ with $\widetilde{\mathbf{w}}^{(i)}$ using the same method for calculating $\widetilde{\mathbf{u}}^{(i)}$.
For SDIRK, the upper bound of the summation in (49d) is $i$. Substitute (61) into (49c,d), separate the terms with and without $r^{(i)}$, switch the summation order by $\sum_{j=1}^{i}a_{i,j}\sum_{\ell=1}^{j}r^{(\ell)}\Delta\mathbf{w}_{\ell}^{(j)}=\sum_{\ell=1}^{i}r^{(\ell)}\sum_{j=\ell}^{i}a_{i,j}\Delta\mathbf{w}_{\ell}^{(j)}$, and we have (GES.3). Then (GES.4) follows from (49g).
(GES.5) follows from substituting the second equation in (61) into (49f),
$\begin{array}[]{l}r^{(i)}=r^{n}+\gamma kI_{cv}\left(\widetilde{\mathbf{u}}^{(i)},\ \overline{\mathbf{u}}^{(i)}+r^{(i)}\mathbf{u}_{i}^{(i)}\right)+k\sum_{j=1}^{i-1}a_{i,j}I_{cv}\left(\widetilde{\mathbf{u}}^{(j)},\ {\mathbf{u}}^{(j)}\right),\end{array}$
and moving the $r^{(i)}$-term on the RHS to the LHS. Finally, the parenthesis on the LHS of (GES.5) must be no less than 1 because of (46), Lemma 3.10, and the inner product of $\mathbf{u}_{i}^{(i)}$ to the equation on $\mathbf{w}_{i}^{(i)}$ in (GES.3).
(48) and (GES.1–6) are the complete algorithmic steps of the GES scheme. Theorems 5.2 and 5.4 state that both the velocity divergence and the modified energy in GES decrease monotonically provided that the employed SDIRK is algebraically stable; an example is given in (62).
In this work, we discretize the continuous spatial operators in Definition 5.5 by the fourth-order collocated finite-volume operators in (zhang16:_GePUP, , Sec. 3 & 4) to obtain a fully discrete GES scheme. However, it is emphasized that what we have proposed is not a single scheme but a space of solvers, each of which can be easily constructed by making menu choices for ‘orthogonal’ policies that span the solver space. The orthogonal structure of this solver space is very conducive to reusing in a black-box manner the legacy of classical finite volume/difference methods and the wealth of theory and algorithms for numerically solving ODEs; see Table 1. It is the GePUP-E formulation that makes this black-box reuse possible.
It would be ideal if the conclusions of Theorems 5.2 and 5.4 could also hold in the fully discrete case. Unfortunately, for the finite volume discretization on collocated grids, the fully discrete counterparts of Theorems 5.2 and 5.4 only hold asymptotically, i.e., in the limit of $k$ and $h$ simultaneously approaching zero. We defer to future research the investigation of suitable spatial discretizations so that Theorems 5.2 and 5.4 also hold in the fully discrete case.
6 Tests
In this section, we test a fully discrete GES scheme by several numerical experiments to confirm the analytic results in previous sections. We employ the fourth-order finite-volume discretizations in zhang16:_GePUP and a fourth-order, algebraically stable SDIRK method proposed by Du, Ju & Lu du19:sdirk ,
$\begin{tabular}[h]{c|ccc}$\gamma$&$\gamma$&&\\ $\frac{1}{2}$&$\frac{1}{2}-\gamma$&$\gamma$&\\ $1-\gamma$&$2\gamma$&$1-4\gamma$&$\gamma$\\ \hline\cr&$\mu$&$1-2\mu$&$\mu$\end{tabular},\qquad\left\{\begin{array}[]{l}\gamma=\frac{1}{\sqrt{3}}\cos\frac{\pi}{18}+\frac{1}{2},\\ \mu=\frac{1}{6(2\gamma-1)^{2}}.\end{array}\right$. (62)
Table 1: Choices of the main orthogonal policies that constitute the tested GES scheme as a fully discrete GePUP-ES-RK.
Orthogonal policies For the tested GES scheme Other possible options
time integration the SDIRK method (62) algebraically stable RK methods
temporal accuracy the fourth order the second, third, and fifth orders
spatial discretization finite volume finite difference
spatial accuracy the fourth order the second order
variable location collocated (as in zhang16:_GePUP ) staggered
estimating $\widetilde{\mathbf{u}}^{(i)}$ and $\widetilde{\mathbf{w}}^{(i)}$ interpolate
in Definitions 5.1 and 5.5 (50) & (EAV-1,2) both in time and in space
Along with other possible options of the major orthogonal policies, we show in Table 1 our choices that constitute the particular GES scheme to be tested. According to the analysis in Section 5, this particular INSE solver should be fourth-order accurate both in time and in space, preserve monotonic decrease of the total kinetic energy, and have the velocity divergence well under control.
All numerical tests are performed on a rectangular domain in Cartesian coordinates, where the electric boundary conditions in (32b) are enforced by homogeneous Dirichlet conditions for the tangential velocity components and a homogeneous Neumann condition for the normal velocity. The initial cell-averaged velocity is calculated by Boole’s rule, a sixth-order formula of Newton-Cotes quadrature. Since exact solutions are unavailable, we define the computational error via Richardson extrapolation, i.e., by the difference of the solution on the current grid and that of the next finer grid.
For different values of the penalty parameter $\lambda=$1, 10, and 100, the corresponding error norms in all numerical tests remain the same up to the first two significant digits; these results differ from those reported in shirokoff11:_navier_stokes ; rosales21:_high_poiss_navier_stokes . However, this is not surprising since projecting $\mathbf{w}$ to $\mathbf{u}$ with $\mathbf{u}\cdot\mathbf{n}=0$ in (32c,d) and setting $\mathbf{w}^{0}=\mathbf{u}^{0}$ in (33) already imply a small magnitude of $\mathbf{n}\cdot\mathbf{w}$. Nonetheless, a positive-valued $\lambda$ guarantees the no-penetration condition be fulfilled. Hereafter we only show results in the case of $\lambda=1$.
6.1 Single-vortex tests
(a) tt=40
(b) tt=60
Figure 2: Snapshots of vorticity for the single-vortex test with $\mathrm{Re}=20,000$ and $\lambda=1$ on a uniform grid with $h=\frac{1}{1024}$ and $\mathrm{Cr}=0.5$. The region of each cell is filled by a single color that corresponds to the cell-averaged vorticity. No image smoothing is applied.
Following bch91:_second_order_projection_for_viscous_incompressible_flow , we define an axisymmetric velocity field on $\Omega=[0,1]^{2}$ by
$u_{\theta}(r_{v})=\left\{\begin{array}[]{cc}\Gamma(\frac{1}{2}r_{v}-4r_{v}^{3})&\text{ if }r_{v}<R;\\ \Gamma\frac{R}{r_{v}}(\frac{1}{2}R-4R^{3})&\text{ if }r_{v}\geq R,\end{array}\right$. (63)
where $r_{v}$ is the distance from the domain center $(\frac{1}{2},\frac{1}{2})^{T}$. The choices $R=0.2$ and $\Gamma=1$ give $\max(u_{\theta})=0.068$. A small viscosity $\nu=3.4\times 10^{-6}$ yields a high Reynolds number $\mathrm{Re}=20,000$. The initial velocity is obtained by projecting cell averages of $\mathbf{u}$ in (63) ten times to make it approximately divergence-free.
Table 2: Errors and convergence rates of the GES scheme in Table 1 for solving the single-vortex test with Re = 20,000, $t_{0}=0.0$, $t_{e}=60$, $\lambda=1$ and the Courant number $\mathrm{Cr}=0.5$.
$h$ $\frac{1}{256}-\frac{1}{512}$ Rate $\frac{1}{512}-\frac{1}{1024}$ Rate $\frac{1}{1024}-\frac{1}{2048}$
$\mathbf{u}$ $L_{\infty}$ 9.44e-04 3.63 7.61e-05 3.86 5.24e-06
$L_{1}$ 3.87e-05 3.76 2.86e-06 3.90 1.91e-07
$L_{2}$ 7.34e-05 3.73 5.52e-06 3.89 3.73e-07
$\nabla\cdot\mathbf{u}$ $L_{\infty}$ 1.08e-02 2.03 2.64e-03 2.82 3.75e-04
$L_{1}$ 5.48e-05 3.38 5.27e-06 3.98 3.35e-07
$L_{2}$ 3.26e-04 2.47 5.87e-05 3.28 6.04e-06
$q$ $L_{\infty}$ 8.96e-06 2.63 1.45e-06 2.94 1.89e-07
$L_{1}$ 7.91e-07 2.41 1.49e-07 2.50 2.63e-08
$L_{2}$ 1.38e-06 2.71 2.10e-07 2.59 3.50e-08
$\nabla q$ $L_{\infty}$ 4.60e-04 2.58 7.67e-05 3.17 8.51e-06
$L_{1}$ 1.78e-05 3.28 1.83e-06 2.89 2.48e-07
$L_{2}$ 3.75e-05 3.24 3.98e-06 3.00 4.96e-07
(a) $\mathcal{E}_{h}^{n}$ as in (65)
(b) $|r_{h}^{n}-1|$
(c) $\left\|\mathbf{D}\mathbf{W}_{h}^{n}\right\|_{\mathcal{C}}$ as in (64)
(d) Zoom-in of (c)
Figure 3: Evolutions of the modified energy, the SAV, and the $L_{2}$-norm of $\mathbf{D}\mathbf{W}_{h}^{n}$ produced by the GES scheme in Table 1 for solving the single-vortex test with $\mathrm{Re}=2\times 10^{4}$, $\mathrm{Cr}=0.5$, and $\lambda=1$. The abscissa in all subplots is time. In (a) and (b), $h=\frac{1}{1024}$ and $k\approx 7.18\times 10^{-3}$. In (c) and (d), the curves marked by $\blacksquare$,” $\bullet$,” and $\blacklozenge$” represent the results for $h=\frac{1}{512}$, $\frac{1}{1024}$, and $\frac{1}{2048}$, respectively.
The tests are performed on four successively refined grids with uniform grid size $h$. The time span $[0,60]$ is made long enough for the turbulent boundary layers to develop prominent Lagrangian coherent structures. The uniform time step size $k$ is given by setting the Courant number $\mathrm{Cr}:=\frac{k}{h}\max(u_{\theta})$ at 0.5. Snapshots of the vorticity at time $t=40$ and at the final time $t=60$ are shown in Figure 2, where the essential features of vortex sheet roll-up and counter-vortices agree with those in bch91:_second_order_projection_for_viscous_incompressible_flow .
It is clear in Table 2 that convergence rates of the velocity are close to 4 in all norms while those of the scalar $q$ and its gradient show order reductions, which are caused by the fact that the Neumann boundary condition in (47g) has to be obtained from spatial derivatives of the velocity and calculating these derivatives incurs order reductions in finite-volume discretizations.
Let $\mathbf{U}_{h}^{n}$ and $\mathbf{W}_{h}^{n}$ denote finite-volume solutions that approximate cell averages of $\mathbf{u}$ and $\mathbf{w}$ at time $t^{n}$, respectively. Then the $L_{2}$-norm for a finite-volume solution $\mathbf{V}_{h}^{n}$ is defined as
$\begin{array}[]{l}\|\mathbf{V}^{n}_{h}\|_{\mathcal{C}}:=\sqrt{\sum_{\mathcal{C}_{\mathbf{i}}}\|\mathcal{C}_{\mathbf{i}}\|\cdot|\mathbf{V}^{n}_{h,\mathbf{i}}|^{2}}\end{array}$ (64)
where $\mathcal{C}_{\mathbf{i}}$ ranges over all control volumes. The modified energy is then
$\begin{array}[]{l}\mathcal{E}_{h}^{n}:=\frac{1}{2}\left(\|\mathbf{U}^{n}_{h}\|_{\mathcal{C}}^{2}+|r_{h}^{n}|^{2}\right),\end{array}$ (65)
where $r_{h}^{n}$ is the computed value of the SAV $r$ at time $t^{n}$.
As shown in Figure 3(a,b), over the entire simulation $\mathcal{E}_{h}^{n}$ decreases monotonically and $|r_{h}^{n}-1|$ remains below $2.0\times 10^{-11}$, indicating that the (unmodified) kinetic energy $\frac{1}{2}\|\mathbf{U}^{n}_{h}\|_{\mathcal{C}}^{2}$ also decreases monotonically.
In Figure 3(c,d), the $L_{2}$-norm of velocity divergence on the coarsest grid $h=\frac{1}{512}$ first decreases dramatically during the first several seconds, then gradually increases, and oscillates within a certain range. Furthermore, the range of oscillation decreases quickly as the grid is refined. To understand this evolution pattern, we consider the fully discrete counterpart of (51), i.e.,
$\left\|\mathbf{D}\mathbf{W}_{h}^{n+1}\right\|_{\mathcal{C}}^{2}-\left\|\mathbf{D}\mathbf{W}_{h}^{n}\right\|_{\mathcal{C}}^{2}\leq-2k\nu\sum_{i=1}^{s}b_{i}\left\|\mathbf{G}\mathbf{D}\mathbf{W}_{h}^{(i)}\right\|_{\mathcal{C}}^{2}+O(h^{p}),$ (66)
where $\mathbf{D}$ is the discrete divergence, $\mathbf{G}$ the discrete gradient, and $O(h^{p})$ errors of spatial discretization. Numerical results such as those in Table 2 show $p>1$.
The key difference between (66) and (51) is the extra term $O(h^{p})$ in (66), which explains why the discrete divergence does not decay monotonically. Since the initial condition of $\mathbf{w}(t_{0})$ for this test is not divergence free, the RHS of (66) is dominated, in the early simulation stage, by the first term that contains $\|\mathbf{G}\mathbf{D}\mathbf{W}^{(i)}_{h}\|_{\mathcal{C}}$. Hence (66) dictates the decay of the discrete divergence. However, as $\|\mathbf{D}\mathbf{W}^{(i)}_{h}\|_{\mathcal{C}}$ decreases continuously, the RHS of (66) eventually becomes dominated by $O(h^{p})$. Then the inequality (66) loses control over the discrete divergence since $O(h^{p})$ is not negative-definite. When the discrete divergence increases to the degree such that the magnitude of $O(h^{p})$ is less than that of the other RHS term, the above pattern is repeated, leading to the oscillation of the discrete divergence. The bottom line is, however, that the discrete divergence is indirectly controlled by the term $O(h^{p})$ and thus the oscillation becomes less prominent as the grid is refined.
6.2 Viscous-box tests
Following bell89:_secon_order_projec_method_incom , we set the initial velocity on $\Omega=[0,1]^{2}$ to
$\mathbf{u}_{0}(x,y)=\left(\begin{array}[]{c}\,\sin^{2}(\pi x)\,\sin(2\pi y)\\ \,-\sin(2\pi x)\,\sin^{2}(\pi y)\end{array}\right)$ (67)
and advance cell-averaged initial values from $t_{0}=0$ to $t_{e}=0.5$ on four successively refined uniform grids. The Courant number is defined as $\mathrm{Cr}:=\frac{k}{h}\|\mathbf{u}_{0}\|_{\infty}$, where $\|\mathbf{u}_{0}\|_{\infty}$ is the max-norm of the initial velocity $\mathbf{u}_{0}$.
Table 3: Errors and convergence rates of the GES scheme in Table 1 for solving the 2D viscous-box test with Re $10^{4}$, $t_{0}=0.0$, $t_{e}=0.5$, $\mathrm{Cr}=0.5$ and $\lambda=1$.
$h$ $\frac{1}{128}-\frac{1}{256}$ Rate $\frac{1}{256}-\frac{1}{512}$ Rate $\frac{1}{512}-\frac{1}{1024}$
$\mathbf{u}$ $L_{\infty}$ 1.02e-03 3.20 1.10e-04 4.09 6.49e-06
$L_{1}$ 4.10e-05 3.53 3.55e-06 3.94 2.32e-07
$L_{2}$ 9.85e-05 3.50 8.71e-06 3.96 5.59e-07
$\nabla\cdot\mathbf{u}$ $L_{\infty}$ 7.33e-03 1.32 2.94e-03 2.44 5.41e-04
$L_{1}$ 3.15e-04 3.18 3.47e-05 3.70 2.68e-06
$L_{2}$ 8.79e-04 2.22 1.89e-04 2.99 2.38e-05
$q$ $L_{\infty}$ 4.17e-04 1.75 1.24e-04 1.96 3.18e-05
$L_{1}$ 6.61e-05 1.23 2.81e-05 1.70 8.69e-06
$L_{2}$ 9.92e-05 1.39 3.78e-05 1.73 1.14e-05
$\nabla q$ $L_{\infty}$ 5.01e-03 1.70 1.54e-03 1.86 4.27e-04
$L_{1}$ 6.82e-04 1.71 2.08e-04 1.86 5.75e-05
$L_{2}$ 1.02e-03 1.69 3.15e-04 1.90 8.43e-05
Table 4: Errors and convergence rates of the GES scheme in Table 1 for solving the 2D viscous-box test with $\text{Re}=10^{2}$, $t_{0}=0.0$, $t_{e}=0.5$, $\mathrm{Cr}=0.1$ and $\lambda=1$.
$h$ $\frac{1}{64}-\frac{1}{128}$ Rate $\frac{1}{128}-\frac{1}{256}$ Rate $\frac{1}{256}-\frac{1}{512}$
$\mathbf{u}$ $L_{\infty}$ 7.84e-06 2.46 1.43e-06 2.78 2.08e-07
$L_{1}$ 2.03e-06 3.99 1.28e-07 4.04 7.76e-09
$L_{2}$ 2.63e-06 3.96 1.69e-07 3.99 1.07e-08
$\nabla\cdot\mathbf{u}$ $L_{\infty}$ 3.90e-04 1.91 1.03e-04 2.02 2.55e-05
$L_{1}$ 1.55e-05 3.83 1.09e-06 3.75 8.11e-08
$L_{2}$ 5.62e-05 3.24 5.96e-06 2.96 7.64e-07
$q$ $L_{\infty}$ 2.37e-04 1.83 6.67e-05 1.91 1.78e-05
$L_{1}$ 3.33e-05 2.04 8.10e-06 2.05 1.95e-06
$L_{2}$ 5.09e-05 2.07 1.21e-05 2.06 2.92e-06
$\nabla q$ $L_{\infty}$ 4.92e-03 0.93 2.57e-03 0.42 1.92e-03
$L_{1}$ 3.24e-04 1.97 8.26e-05 2.01 2.05e-05
$L_{2}$ 5.58e-04 1.72 1.69e-04 1.82 4.79e-05
(a) $\mathcal{E}_{h}^{n}$ for Re $10^{2}$ and $\mathrm{Cr}=0.1$
(b) $\mathcal{E}_{h}^{n}$ for Re $10^{4}$ and $\mathrm{Cr}=0.5$
(c) $|r_{h}^{n}-1|$ for Re $10^{2}$ and $\mathrm{Cr}=0.1$
(d) $|r_{h}^{n}-1|$ for $\mathrm{Re}=10^{4}$ and $\mathrm{Cr}=0.5$
(e) $\left\|\mathbf{D}\mathbf{W}_{h}^{n}\right\|_{\mathcal{C}}$ (solid markers) and $\left\|\mathbf{D}\mathbf{U}_{h}^{n}\right\|_{\mathcal{C}}$ (hollow markers) for Re $10^{2}$ and $\mathrm{Cr}=0.1$
(f) $\left\|\mathbf{D}\mathbf{W}_{h}^{n}\right\|_{\mathcal{C}}$ (solid markers) and $\left\|\mathbf{D}\mathbf{U}_{h}^{n}\right\|_{\mathcal{C}}$ (hollow markers) for Re $10^{4}$ and $\mathrm{Cr}=0.5$
Figure 4: Results of the GES scheme in Table 1 for solving the 2D viscous-box test with $h=\frac{1}{512}$ and $\lambda=1$; see (64) and (65) for precise definitions of $\left\|\cdot\right\|_{\mathcal{C}}$ and $\mathcal{E}_{h}^{n}$. The abscissa in all subplots is time.
Errors and convergence rates in the cases of $\mathrm{Re}=10^{4}$ and $\mathrm{Re}=10^{2}$ are shown in Tables 3 and 4, respectively. For $\mathrm{Re}=10^{4}$, convergence rates of the velocity in all norms are close to four. In contrast, those for $\mathrm{Re}=10^{2}$ are close to four in the $L_{1}$-norm and the $L_{2}$-norm, but are around 2.5 in the $L_{\infty}$-norm. Accordingly, convergence rates of $\nabla q$ in the $L_{\infty}$-norm for $\mathrm{Re}=10^{4}$ are also substantially higher than those for $\mathrm{Re}=10^{2}$.
These results are not out of expectations. We have proved in Theorem 3.7 that the pressure $q$ converges to the pressure $p$ and have shown in Section 2.3 that the pressure gradient $\nabla p$ in the INSE can be split into two parts $\nabla p=\nabla p_{c}+\nu\nabla p_{s}$, where $\nabla p_{s}$ responses to the Laplace-Leray commutator. When $\nu$ is sufficiently large, $\nu\nabla p_{s}$ dominates $\nabla p_{c}$ and accounts for the bulk of $\nabla p$. Cozzi and Pego cozzi11:_laplac_leray showed that $\|\nabla p_{s}\|$ may not be bounded at a boundary point that is not $C^{3}$. As a practical interpretation, the pressure could develop steep gradient at a $C^{1}$ discontinuity of the domain boundary for low-Reynolds-number flows. Therefore, we believe that the order reduction in the case of Re $10^{2}$ is caused by the dominance of $\nu\nabla p_{s}$ and the sharp corners ($C^{1}$ discontinuities) of the square domain.
As shown in Figure 4(a,b), the modified energy $\mathcal{E}_{h}^{n}$ decreases monotonically over the entire simulation for both $\mathrm{Re}=10^{4}$ and $\mathrm{Re}=10^{2}$. It is also clear that $\mathcal{E}_{h}^{n}$ decreases faster in the higher-viscosity case of Re $10^{2}$, which confirms Theorem 5.4. Figure 4(c,d) show that the deviation of SAV from 1 is at most $10^{-10}$, thus the (unmodified) kinetic energy also decreases monotonically.
For both $\mathrm{Re}=10^{4}$ and $\mathrm{Re}=10^{2}$, the evolution of the $L_{2}$-norm of velocity divergence shown in Figure 4(e,f) has essentially the same pattern: the $L_{2}$-norm first increases to a local maximum and then decreases. This pattern is different from that shown in Figure 3(c,d), but can still be very well explained by (66). Since the initial velocity (67) is divergence-free, the magnitude of $\|\mathbf{G}\mathbf{D}\mathbf{W}^{(i)}_{h}\|_{\mathcal{C}}$ is small in the early simulation stage, during which the inequality (66) has no control over the discrete velocity divergence yet. However, as the discrete divergence accumulates to the point when the RHS of (66) gets dominated by its first term, the inequality (66) takes effect and forces the discrete divergence to decrease.
6.3 Viscous-box tests with initially non-solenoidal velocity
Tests in this subsection are the same as those in Section 6.2 except that the initial condition is set to the following non-solenoidal velocity,
$\begin{array}[]{l}\mathbf{w}_{0}=\mathbf{u}_{0}+\epsilon\nabla\phi,\quad\text{ with }\phi=\frac{1}{2\pi^{2}}\sin^{2}(\pi x)\sin^{2}(\pi y),\end{array}$ (68)
where $\mathbf{u}_{0}$ is the initially solenoidal velocity in (67), $\epsilon\nabla\phi$ the perturbation of $\mathbf{u}_{0}$ away from the divergence-free space, and $\epsilon>0$ the parameter to control the magnitude of the perturbation.
(a) $\epsilon=1$, $t=t_{0}=0$
(b) $\epsilon=0$, $t=t_{0}=0$
(c) $\epsilon\!=\!10^{-3}$, $t=t_{0}=0$
(d) $\epsilon=1$, $t=0.5$, Re $10^{4}$
(e) $\epsilon=0$, $t=0.5$, Re $10^{4}$
(f) $\epsilon\!=\!10^{-3}$, t=0.5t\!=\!0.5, Re $10^{4}$
(g) $\epsilon=1$, $t=0.5$, Re $10^{2}$
(h) $\epsilon=0$, $t=0.5$, Re $10^{2}$
(i) $\epsilon\!=\!10^{-3}$, $t=0.5$, Re $10^{2}$
[…]
Hereafter we denote by $\mathbf{U}^{n}_{\epsilon=0}$ the reference solution obtained from the divergence-free initial velocity $\mathbf{u}_{0}$, i.e., $\mathbf{U}^{n}_{h}$ in Section 6.2, and denote by $\mathbf{U}^{n}_{\epsilon>0}$ the perturbed solution computed from a perturbed initial velocity $\mathbf{w}_{0}$.
Streamlines of the initial velocity and the final solutions are plotted in Figure 5, where no qualitative differences can be observed between corresponding plots in the third and the second columns, due to the small perturbation magnitude $\epsilon=10^{-3}$. In contrast, for $\epsilon=1$, the perturbation is sufficient to generate different patterns for the initial velocity, cf. Figure 5(a,b). For example, the streamlines in Figure 5(a) are not closed due to the large magnitude of velocity divergence. However, the perturbation effects have already died out at $t=0.5$ and the final perturbed solution $\mathbf{U}^{n}_{\epsilon=1}$ is visually indistinguishable from the unperturbed solution $\mathbf{U}^{n}_{\epsilon=0}$ for both Re $10^{4}$ and Re $10^{2}$; see Figure 5(d,e) and Figure 5(g,h).
How fast does the perturbed solution $\mathbf{U}^{n}_{\epsilon>0}$ converge to the solution $\mathbf{U}^{n}_{\epsilon=0}$? The answer lies in Figure 6, where the deviation of $\mathbf{U}^{n}_{\epsilon>0}$ from $\mathbf{U}^{n}_{\epsilon=0}$ decays by a factor of at least $10^{3}$ within the first time step for all test cases. Furthermore, although the norm $\|\mathbf{U}^{n}_{\epsilon>0}-\mathbf{U}^{n}_{\epsilon=0}\|_{\mathcal{C}}$ appears to be proportional to the perturbation magnitude, the reduction rate during the first time step is largely independent of the time step size; see Figure 6(b,c) and Figure 6(e,f).
(a) Re $10^{2}$, Cr =0.1=0.1
(b) Re $10^{2}$, Cr =0.1=0.1
(c) Re $10^{2}$, Cr =0.01=0.01
(d) Re $10^{4}$, Cr =0.5=0.5
(e) Re $10^{4}$, Cr =0.5=0.5
(f) Re $10^{4}$, Cr =0.05=0.05
Figure 6: Evolutions of $\|\mathbf{U}^{n}_{\epsilon>0}-\mathbf{U}^{n}_{\epsilon=0}\|_{\mathcal{C}}$, i.e., the $L_{2}$-norm (64) of the difference between the computed velocities of the perturbed viscous-box tests in Section 6.3 on a uniform spatial grid $h=\frac{1}{512}$. The markers $\bullet$,” $\blacksquare$,” $\blacktriangle$” and “++” represent values of $\|\mathbf{U}^{n}_{\epsilon>0}-\mathbf{U}^{n}_{\epsilon=0}\|_{\mathcal{C}}$ for $\epsilon=1$, $0.1$, $10^{-2}$, and $10^{-3}$, respectively. The abscissas in subplots (b,c,e,f) and (a,d) represent the index of time steps and the simultation time, respectively.
(a) $\mathcal{E}^{n}_{h}$ for Re $10^{2}$ and Cr = 0.1
(b) $\mathcal{E}^{n}_{h}$ for Re $10^{4}$ and Cr = 0.5
(c) $|r^{n}_{h}-1|$ for Re $10^{2}$ and Cr = 0.1
(d) $|r^{n}_{h}\!-\!1|$ for Re $10^{4}$ and Cr = 0.5
(e) $\|\mathbf{DU}^{n}_{\epsilon}\|_{\mathcal{C}}$ for Re $10^{2}$ and Cr = 0.1
(f) $\|\mathbf{DU}^{n}_{\epsilon}\|_{\mathcal{C}}$ for Re $10^{4}$ and Cr = 0.5
Figure 7: Evolutions of $\mathcal{E}^{n}_{h}$, $|r^{n}_{h}-1|$, and $\|\mathbf{DU}^{n}_{\epsilon}\|_{\mathcal{C}}$ produced by the GES scheme in Table 1 for solving the perturbed viscous box tests on a uniform grid of $h=\frac{1}{512}$; see (64) and (65) for precise definitions of $\left\|\cdot\right\|_{\mathcal{C}}$ and $\mathcal{E}_{h}^{n}$. The dashed curves represent results with $\epsilon=0$ (i.e., those in Figure 4) while those marked by $\bullet$,” $\blacksquare$,” $\blacktriangle$,” and $\bigstar$” correspond to $\epsilon=1$, $10^{-1}$, $10^{-2}$ and $10^{-3}$, respectively. The abscissa of all subplots is time.
For two curves very close to each other, their tangent vectors might differ largely, so do other geometric quantities such as the curvature. Likewise, $\|\mathbf{U}^{n}_{\epsilon>0}-\mathbf{U}^{n}_{\epsilon=0}\|_{\mathcal{C}}$ being small does not imply $\|{\mathbf{D}}\mathbf{U}^{n}_{\epsilon>0}-{\mathbf{D}}\mathbf{U}^{n}_{\epsilon=0}\|_{\mathcal{C}}$ being small. This issue is addressed by Figure 7, which shows the perturbation effects on evolutions of the modified energy, the SAV, and the velocity divergence.
Compared to Figure 4(a,b), Figure 7(a,b) feature a decrease of the modified energy during the very first time step for $\epsilon=1$, indicating that most of the extra energy corresponding to the non-solenoidal perturbation has been removed. Values of $|r_{h}^{n}-1|$ in Figure 7(c,d) are much larger than those in Figure 4(c,d), which is still acceptable because the deviation of $r_{h}^{n}$ from 1 is about $10^{-5}$ even for the case of the largest perburbation $\epsilon=1$. In Figure 7(e,f), the evolution curve of $\|\mathbf{DU}^{n}_{\epsilon}\|_{\mathcal{C}}$ for each $\epsilon>0$ overlaps with that of $\epsilon=0$ except for the first four steps.
(a) Re $10^{2}$ and Cr = 0.1
(b) Re $10^{4}$ and Cr = 0.5
(c) Re $10^{2}$ and Cr = 0.01
(d) Re $10^{4}$ and Cr = 0.05
Figure 8: Evolution of discrete velocity divergence produced by the GES scheme in Table 1 for solving the perturbed viscous box tests on a uniform grid of $h=\frac{1}{512}$. The abscissa of all subplots is the index of time steps. Subplots (a,b) can be considered as more detailed versions of Figure 7 (e,f). Values of $\|\mathbf{DU}^{n}_{\epsilon}\|_{\mathcal{C}}$ for $\epsilon=0,10^{-3},1$ are represented respectively by solid markers $\blacktriangle$,” $\blacksquare$,” and $\bullet$” while $\|\mathbf{DW}^{n}_{\epsilon}\|_{\mathcal{C}}$ respectively by the corresponding the hollow markers $\triangle$,” $\square$,” and $\circ$.” Results of $\epsilon=10^{-1},10^{-2}$ are in between those of $\epsilon=10^{-3},1$ and have the same qualitative pattern.
To examine what happens in these four steps, the temporal variations of the $L_{2}$-norm of the velocity divergence are plotted in Figure 8 for the two representative cases of $\epsilon=1,10^{-3}$. Different from previous figures on velocity divergence, Figure 8 contains values of both $\|\mathbf{DW}^{n}_{\epsilon}\|_{\mathcal{C}}$ and $\|\mathbf{DU}^{n}_{\epsilon}\|_{\mathcal{C}}$, which are connected by the discrete projection that approximates $\mathscr{P}$} in (48). As illustrated in Figure 8, the discretization error might increase $\|\mathbf{DW}^{n}_{\epsilon}\|_{\mathcal{C}}$ during a single time step, but the magnitude of this increase approaches a constant value after about eight time steps; furthermore, the discrete projection at the end of a time step counteracts this increase. In each subplot of Figure 8, the sequence of solid markers $\bullet$” and $\blacksquare$” during the first four time steps clearly demonstrates the fast decay of $\|\mathbf{DU}^{n}_{\epsilon}\|_{\mathcal{C}}$. In addition, a comparison of the two rows of Figure 8 shows that the adverse effects of the perturbation upon the divergence-free condition can be further reduced by decreasing the size of the initial time steps.
We have demonstrated two aspects of the prominent advantage of GePUP-ES in handling an initially non-solenoidal velocity. First, the velocity divergence decays exponentially. Second, Figures 6 and 8 suggest that the time period of adverse effects caused by the initially non-solenoidal velocity can be very much shortened by reducing the size of the first several time steps. We emphasize that theoretically there is no guarantee of the decay of velocity divergence for any initially non-solenoidal velocity yet, as the well-posedness of INSE is still one of the unsolved Millennium Prize Problems.
(a) $\mathcal{E}^{n}_{h}$
(b) $\|\mathbf{DU}^{n}_{\epsilon}\|_{\mathcal{C}}$
Figure 9: Evolution of $\mathcal{E}^{n}_{h}$ and $\|\mathbf{DU}^{n}_{\epsilon}\|_{\mathcal{C}}$ produced by applying the MOL approach to the strong form of UPPE in (13) for Re = $10^{4}$ , Cr = 0.5, and $h=\frac{1}{512}$. The curves marked by $\bullet$” and $\blacksquare$” correspond to $\epsilon=1$ and $10^{-3}$, respectively, while the dashed curve corresponds to $\epsilon=0$. The abscissa in all subplots is time.
The finite-volume based MOL scheme is also applied to the strong form of UPPE in (13) to obtain solutions of the perturbed viscous-box tests. As shown in Figure 9, both the velocity divergence and the total kinetic energy blow up, even for the case of an initially solenoidal velocity. It is also clear that the blow-up time becomes sooner as $\epsilon$ gets larger.
Results of viscous-box tests solved by GePUP, as shown in (zhang16:_GePUP, , Table 3), are very close to those in Table 3 for $\mathrm{Re}=10^{4}$. In the case of $\mathrm{Re}=10^{2}$, the error norms of $\mathbf{u}$ produced by GePUP-ES on the finest grid are approximately 20% smaller than those by GePUP, cf. Table 4 and (zhang16:_GePUP, , Table 4). For the perturbed viscous box tests in this subsection, results of GePUP also show fast decays of discrete velocity divergence. However, GePUP-ES is a decisive advance from GePUP in that, in the semi-discrete case, the decay of velocity divergence and kinetic energy can be rigorously proven.
7 Conclusions
We have shown that the INSE with no-slip conditions can be equivalently reformulated as variants of the GePUP formulation zhang16:_GePUP , where the main evolutionary variable is a non-solenoidal velocity with electric boundary conditions whose divergence, controlled by a heat equation with homogeneous Dirichlet boundary conditions, decays exponentially. This GePUP-E reformulation is suitable for numerically solving the INSE because
  • time integration and spatial discretization are completely decoupled so that high-order INSE solvers can be easily obtained from menu choices of orthogonal policies,
  • the constituting modules such as a time integrator are employed in a black-box manner so that no internal details of any module are needed in building the INSE solver,
  • the influences of nonzero velocity divergence upon numerical stability and accuracy are clear,
  • a coupling of GePUP-E to SAV yields semi-discrete schemes with monotonically decreasing kinetic energy.
Results of numerical experiments confirm the analysis.
We are currently augmenting the GePUP-ES solver to parallel computing and adaptive mesh refinement zhao2025FourthOrderAMR for an enhanced resolution of the multiple time scales and length scales in flows at moderate or high Reynolds numbers. Another work in progress is the development of GePUP solvers for the INSE with irregular and moving boundaries via poised lattice generation zhang2024PLG .
The next step along this line of research is the design of fully discrete algorithms that ensure decays of velocity divergence and total kinetic energy. We also plan to consider other types of boundary conditions such as the nonhomogeneous Dirichlet conditions, the radiation conditions, and mixed conditions.
Acknowledgements.
This work was supported by the grant 12272346 from the National Natural Science Foundation of China. We acknowledge helpful discussions with Prof. Buyang Li at the Hong Kong Polytechic University. We are also grateful to one anonymous referee, whose insightful comments and suggestions lead to an improvement of this paper.
Declarations
Data Availability
The datasets generated during and/or analysed during the current study are available from the corresponding author on reasonable request.
Conflict of interest
The authors declare that they have no conflict of interest.